### v1

# Benign Overfitting for $\alpha$ Sub-exponential Input

Abstract
This paper investigates the phenomenon of benign overfitting in binary classification problems with heavy-tailed input distributions. We extend the analysis of maximum margin classifiers to $\alpha$ sub-exponential distributions, where $\alpha\in(0,2]$, generalizing previous work that focused on sub-gaussian inputs. Our main result provides generalization error bounds for linear classifiers trained using gradient descent on unregularized logistic loss in this heavy-tailed setting. We prove that under certain conditions on the dimensionality $p$ and feature vector magnitude $\|\mu\|$, the misclassification error of the maximum margin classifier asymptotically approaches the noise level. This work contributes to the understanding of benign overfitting in more robust distribution settings and demonstrates that the phenomenon persists even with heavier-tailed inputs than previously studied.
1 Introduction
In the field of machine learning, a phenomenon that contradicts the long-standing intuition of statistical learning theory has been garnering attention. This phenomenon is called benign overfitting. According to conventional theory, when a model excessively fits the training data, its generalization performance on unseen data was expected to decline. However, experiments using deep neural networks have revealed that models that perfectly fit noisy training data surprisingly demonstrate good performance on unseen data as well [32, 2].
This phenomenon suggests a significant gap between machine learning theory and practice, attracting the attention of researchers. To deepen our understanding of benign overfitting, studies have been conducted in simpler statistical settings that are more amenable to theoretical analysis, such as linear regression [13, 1, 24, 25, 28, 9, 8], sparse linear regression [15, 7, 16, 29], logistic regression [22, 6, 19, 23, 30, 21, 11, 33], and kernel-based estimators [3, 20, 17, 18]. These studies are rapidly advancing our understanding of the conditions and mechanisms under which benign overfitting occurs.
Particularly in binary classification problems, a standard mixture model setting is often considered [6, 30, 11, 33]. This setting involves the classification of well-separated data under adversarially corrupted labels, where the input distribution is assumed to be sub-gaussian. However, benign overfitting in settings with input distributions heavier than sub-gaussian, that is, settings more robust to input variations, has not been extensively discussed.
Interestingly, numerical experiments suggest that benign overfitting can occur even in standard mixture settings with heavier-tailed distributions than the normal distribution, as illustrated in Figure 1. This observation further emphasizes the need to investigate benign overfitting in a broader range of distributional settings.
Figure 1: Training and Test Error vs. Dimension ($p$) for Different $\gamma$ Values. The plot shows the test error and train error versus the dimension $p$ for a maximum margin classifier. The number of training samples $n_{\mathrm{train}}=200$ is kept constant, and the number of test samples $n_{\mathrm{test}}$ is $1000$. The number of parameters $p$ varies from $100$ to $1500$. The data is generated from a relaxed setting of a standard mixture model; two classes form a symmetric mixture of generalized normal distributions (see Example 2) with location parameter $x_{0}=[1,\cdots,1]$ and $x_{0}=[-1,\cdots,-1]$, shape parameter $\gamma=0.25,0.50,2$, controlling the tail heaviness, and scale parameter $\sigma$ adjusted so that the variance becomes 1 for each $\gamma$. The noise level $\eta$ is set to $0.05$, represented by the dotted line. The solid and dashed lines show the training and test errors respectively for different tail heaviness $0.25$ (blue), $0.5$ (green), and $2$ (red), with error bars indicating variability across $50$ trials. The training error remains near zero across all dimensions, while the test error initially decreases and then stabilizes around the noise level as the dimension increases.
In this work, we consider binary classification where the input distribution has heavier tails than sub-gaussian, specifically $\alpha$ sub-exponential with $\alpha\in(0,2]$. The research aims to prove generalization error bounds demonstrating benign overfitting for a linear classifier using gradient descent on the unregularized logistic loss. This work generalizes the bound of Chatterji and Long [6].
1.1 Related works
Benign overfitting in classification: Most related to our work is the theoretical analysis of benign overfitting in classification settings. This line of research aims to understand why classifiers that perfectly fit noisy training data can still generalize well to unseen data. Wang and Thrampoulidis [31] studied a setting where the two classes are a symmetric mixture of gaussian (or sub-gaussian) distributions, without label noise. Chatterji and Long [6] studied overparameterized linear logistic regression on sub-gaussian mixture models with label flipping noise. They showed how gradient descent can train these models to achieve nearly optimal population risk. Cao et al. [5] extended this work, tightening the upper bound in the case without label flipping noise and establishing a matching lower bound for overparameterized maximum margin interpolators. Wang et al. [30] extended the analysis of maximum margin classifiers to multiclass classification in overparameterized settings. Frei et al. [11] proved that a two-layer fully connected neural network exhibits benign overfitting under certain conditions, such as well-separated log-concave distribution and smooth activation function. Cao et al. [4] focused on the benign overfitting of two-layer convolutional neural networks.
2 Preliminaries
In this section, we introduce the definition of $\alpha$ sub-exponential random variables, the assumptions on the data generation process, and the maximum margin algorithm we consider.
2.1 Notation
In this paper, we use the notation $[n]$ to denote the set $\{1,2,\ldots,n\}$ for a positive integer $n$. For a vector $x$, we use $\|x\|$ to denote its $\ell^{2}$ norm. For a matrix $A$, we use $\|A\|_{\mathrm{HS}}$ to denote its Hilbert–Schmidt norm and $\|A\|_{\mathrm{op}}$ to denote its operator norm.
2.2 $\alpha$ sub-exponential random variable
$\alpha$ sub-exponential random variables are random variables which have exponential type tails.
Definition 1 ($\alpha$ sub-exponential random variable).
A random variable $X$ is called $\alpha$ sub-exponential if there is a positive constant $c_{\alpha}$ such that it holds
$\mathbb{P}\left[|X-\mathbb{E}[X]|\geq t\right]\leq 2\exp\left(-\frac{t^{\alpha}}{c_{\alpha}}\right)$
for any $t>0$. This is equivalent to having a finite exponential Orlicz norm:
$\|X\|_{\psi_{\alpha}}:=\inf\left\{t>0:\mathbb{E}\left[\exp\left(\frac{|X|^{\alpha}}{t^{\alpha}}\right)\right]\leq 2\right\}<\infty.$
If $\alpha=2$, we call the distribution sub-gaussian. If $\alpha=1$, we call it sub-exponential. Here is an example of an $\alpha$ sub-exponential distribution:
Example 2 (Generalized normal distribution).
The probability density function of the generalized normal distribution is defined as:
$f(x;x_{0},\sigma,\gamma)=\frac{\gamma}{2\sigma\Gamma(1/\gamma)}\exp\left(-\left|\frac{x-x_{0}}{\sigma}\right|^{\gamma}\right)$
where $\Gamma$ denotes the gamma function, $x_{0}$ is the location parameter, $\sigma>0$ is the scale parameter, and $\gamma>0$ is the shape parameter. Let $X$ be a random variable following the Generalized Normal Distribution with location parameter $x_{0}=0$. Then $\gamma$ is the maximum value of $\alpha$ such that $\|X\|_{\psi_{\alpha}}$ is finite, and its exponential Orlicz norm is given by
$\|X\|_{\psi_{\gamma}}=\frac{\sigma}{(1-2^{-\gamma})^{1/\gamma}}.$
2.3 Data generation process
We consider a heavy-tailed setting for binary classification, which is a relaxed setting of a standard mixture model setting (Chatterji and Long, 2021 [6]; Frei et al., 2022 [11]). We first define a “clean” distribution $\tilde{P}$ and then define the target distribution $P$ based on $\tilde{P}$:
  1. 1.
    Sample a “clean” label $\tilde{y}\in\{\pm 1\}$ uniformly at random, $\tilde{y}\sim\mathrm{Uniform}(\{\pm 1\})$.
  2. 2.
    Sample $q\sim P_{\mathrm{clust}}$ that satisfies:
    • $P_{\mathrm{clust}}:=P_{\mathrm{clust}}^{(1)}\times\cdots\times P_{\mathrm{clust}}^{(p)}$ is an arbitrary product distribution on $\mathbb{R}^{p}$ whose marginals are all mean-zero with the exponential Orlicz norm at most one, $\|X\|_{\psi_{\alpha}}$ if $X\sim P_{\rm clust}^{(j)}$.
    • For some $\kappa>0$, it holds that $\mathbb{E}_{q\sim P_{\mathrm{clust}}}[\|q\|^{2}]\geq\kappa p$.
  3. 3.
    For an arbitrary orthogonal matrix $U\in\mathbb{R}^{p\times p}$ and $\mu\in\mathbb{R}^{p}$, generate $\tilde{x}=Uq+\tilde{y}\mu$.
  4. 4.
    Let $\tilde{P}$ be the distribution followed by $(\tilde{x},\tilde{y})$.
  5. 5.
    For $\eta\in[0,1]$, let $P$ be an arbitrary distribution on $\mathbb{R}^{p}\times\{\pm 1\}$ that satisfies:
    • All marginal distributions of $P$ are the same as $\tilde{P}$.
    • Total variation between $P$ and $\tilde{P}$ is at most $\eta$.
The reason we assume the $\alpha$ sub-exponential norm of each component is at most $1$ is only for simplifying the proofs and does not affect the main results of the paper, since rescaling the data does not affect the accuracy of the maximum margin algorithm.
This setting is a modification of Chatterji and Long’s framework [6], where we have replaced the sub-gaussian norm with an exponential Orlicz norm. Moreover, it can be observed that when $\alpha=2$, this setting encompasses the original framework.
2.4 Maximum margin algorithm
We consider a linear classifier that takes the form $\mathrm{sign}(\theta\cdot x)$ trained by gradient descent as
$\theta^{(t+1)}=\theta^{(t)}-\beta\nabla R(\theta^{(t)})\quad\mathrm{where}\quad R(\theta):=\sum_{i=1}^{n}\log(1+\exp(-y_{i}\theta\cdot x_{i})),$
where $\beta$ is the learning rate. In Soudry et al., 2018 [27], they prove that if the dataset is linearly separable, in the large-$t$ limit, the normalized parameter of this classifier converges to the hard margin predictor:
$\lim_{t\rightarrow\infty}\frac{\theta^{(t)}}{\|\theta^{(t)}\|}=\frac{w}{\|w\|},$
$w:=\underset{u\in\mathbb{R}^{p}}{\mathrm{argmin}}\|u\|$
$\text{such that }y_{i}(u\cdot x_{i})\geq 1,\text{ for any }i\in[n].$
They have proved this for a certain loss function with a certain smoothness.
2.5 Assumptions
We assume that $\alpha$ and $\kappa$ are fixed constants. We will prove that there exists large enough $C$, depending only on $\alpha$ and $\kappa$, such that our theorem holds for $C$.
We assume the following assumptions:
  1. (A1)
    The failure probability satisfies $\delta\in(0,\frac{1}{C})$,
  2. (A2)
    The number of samples satisfies $n\geq C\log\frac{1}{\delta}$,
  3. (A3)
    The dimension satisfies $p\geq C\max\{\|\mu\|^{2}n,n^{2}\left(\log\frac{n}{\delta}\right)^{\frac{2}{\alpha}}\}$,
  4. (A4)
    The norm of the mean satisfies $\|\mu\|\geq C\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}$,
  5. (A5)
    The learning rate satisfies
    $\beta\leq\frac{1}{2c_{1}}\left(p+2n\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right)^{-1},$
    where $c_{1}=\max\left(\frac{8}{\kappa},\frac{8}{\alpha}\Gamma\left(\frac{2}{\alpha}\right)+\kappa+2\right).$
When $\alpha=2$, our assumptions correspond to that of Chatterji and Long [6].
Example 3 (Generalized noisy rare-weak model).
For any $\alpha\in(0,2]$, the model described above includes a special case called the generalized noisy rare-weak model, which is defined as follows:
  • For any $j\in[p]$, $P_{\mathrm{clust}}^{(j)}$ is a generalized normal distribution with location parameter $x_{0}=0$, shape parameter $\gamma=\alpha$, and scale parameter $\sigma$.
  • The mean vector $\mu\in\mathbb{R}^{p}$ has only $s$ non-zero components, all of which are equal to $\lambda>0$, where $s$ and $\lambda$ are set appropriately to satisfy assumptions (A1)-(A4).
If we require the exponential Orlicz norm to be less than 1, we need to adjust the scale parameter $\sigma$ of $P_{\text{clust}}^{(j)}$. Donoho and Jin [10] studied this model where $\eta=0,\gamma=1$ and $\sigma=1$.
3 Main Results
We derive a generalization bound for the maximum margin classifier in a relaxed standard mixture model.
Theorem 4.
For any $\alpha\in(0,2]$ and $\kappa\in(0,1)$, there exists a constant $c>0$ such that, under assumptions (A1)-(A5), for all large enough C, with probability at least $1-\delta$, the maximum margin classifier $w$ satisfies
$\underset{(x,y)\sim P}{\mathbb{P}}\left[\mathrm{sign}(w\cdot x)\neq y\right]\leq\eta+\exp\left(-c\frac{\|\mu\|^{2\alpha}}{p^{\alpha/2}}\right).$
Building on these results, we can further analyze the implications of Theorem 4. This theorem reveals the relationship between the number of dimensions $p$ and $\|\mu\|$ in determining the success of learning. Specifically, when $\|\mu\|$ increases as $\|\mu\|=\Theta(p^{d})$ for any $d\in(1/4,1/2]$, the misclassification error of the maximum margin classifier asymptotically approaches the noise level $\eta$. This suggests that in high-dimensional feature spaces, if the signal is sufficiently strong, learning can be achieved while minimizing the impact of noise. These findings align with previous research. The rate of increase in $\|\mu\|$ for benign overfitting corresponds to that proved by Chatterji and Long[6] when $\alpha=2$.
Here are the implications of Theorem 4 in the noisy rare-weak model where the mean vector $\mu$ has only $s$ non-zero elements and all non-zero elements equal $\gamma$.
Corollary 5.
There exists a constant $c>0$ such that, under the assumptions (A1)-(A5), in the noisy rare-weak model, for any $\lambda\geq 0$ and all large enough $C$, with probability $1-\delta$, a maximum margin classifier $w$ satisfies
$\underset{(x,y)\sim P}{\mathbb{P}}\left[\mathrm{sign}(w\cdot x)\neq y\right]\leq\eta+\exp\left(-c\frac{(\lambda^{2}s)^{\alpha}}{p^{\alpha/2}}\right).$
Let $\lambda$ be a constant. Jin [14]demonstrated that for the noiseless rare-weak model, learning is impossible when $s=O(\sqrt{p})$ under the Gaussian assumption. Considering the fact that the gaussian distribution is an $\alpha$ sub-gaussian for every $\alpha$ in $(0,2]$, their counterexample can show that our upper bound has optimality in a sense. Strictly speaking, to fit Jin’s model to our model, we need to adjust the scale parameter $\sigma$ of $P_{\text{clust}}^{(j)}$ to make the exponential Orlicz norm less than 1. However, this adjustment does not affect the accuracy of the maximum margin classifier.
4 Proof of Theorem 4
In the Propositions and Lemmas of this section, we assume (A1)-(A5). For simplicity, we assume $U=I$. This assumption can be made without loss of generality for the following reasons:
  • Transformation of the maximum margin classifier:
    If $w$ is the maximum margin classifier for the original data points $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$, then $Uw$ becomes the maximum margin classifier for the transformed data points $(Ux_{1},y_{1}),\ldots,(Ux_{n},y_{n})$.
  • Probability equivalence:
    The probability of misclassification remains unchanged whether we consider $y(w\cdot x)<0$ or $y(Uw)\cdot(Ux)<0$.
For the same reason as in section 4 of [6], without loss of generality, we can assume $\mathbb{P}(x=\tilde{x})=1$ and $\mathbb{P}(y\neq\tilde{y})=\eta$.
Let $\mathcal{S}:=\{(x_{1},y_{1}),\cdots,(x_{n},y_{n})\}$ be samples drawn according to $P$. We define the sets of indices of “noisy” and “clean” samples.
Definition 6.
Let $\mathcal{N}$ denote the set $\{k:y_{k}\neq\tilde{y}_{k}\}$ of indices of “noisy” samples, and $\mathcal{C}$ denote the set $\{k:y_{k}=\tilde{y}_{k}\}$ indices of “clean” samples.
Next, we define $z_{k}$, $\tilde{z}_{k}$, $\xi_{k}$, and $\tilde{\xi}_{k}$ to simplify the subsequent discussion.
Definition 7.
For index $k\in[n]$ of each example, let $z_{k}$ denote $x_{k}y_{k}$ and let $\tilde{z}_{k}$ denote $\tilde{x}_{k}\tilde{y}_{k}$. Let $\xi_{k}$ denote $z_{k}-\mathbb{E}[z_{k}]$ and let $\tilde{\xi}_{k}$ denote $\tilde{z}_{k}-\mathbb{E}[\tilde{z}_{k}]$.
Then, $\xi_{k}$ and $\tilde{\xi}_{k}$ are $\alpha$ sub-exponential, and the following proposition holds:
Proposition 8.
For any $k\in[n]$,
  1. 1.
    $\mathbb{E}[z_{k}]=\mathbb{E}[\tilde{z}_{k}]=\mu$.
  2. 2.
    Each component of $\xi_{k}$ and $\tilde{\xi}_{k}$ is $\alpha$ sub-exponential, with its exponential Orlicz norm at most 1.
Proof.
By definition of $\tilde{z}_{k}$ and $z_{k}$,
$\mathbb{E}[\tilde{z}_{k}]=\mathbb{E}[\tilde{x}_{k}\tilde{y}_{k}]=\mathbb{E}[(q+\tilde{y}_{k}\mu)\tilde{y}_{k}]=\mu,$
$\mathbb{E}[z_{k}]=\mathbb{E}[\tilde{x}_{k}y_{k}]=\mathbb{E}[(q+y_{k}\mu)y_{k}]=\mu.$
Together with the definition of $\xi_{k}$,
$\|\xi_{kl}\|_{\psi_{\alpha}}=\left\|q_{l}y_{k}\right\|_{\psi_{\alpha}}=\left\|q_{l}\right\|_{\psi_{\alpha}}\leq 1,$
$\|\tilde{\xi}_{kl}\|_{\psi_{\alpha}}=\left\|q_{l}\tilde{y}_{k}\right\|_{\psi_{\alpha}}=\left\|q_{l}\right\|_{\psi_{\alpha}}\leq 1.$
The next lemma provides an upper bound for the misclassification error. This bound is expressed in terms of two factors:
  • The expected value of the margin on unperturbed data points, denoted as
    $\underset{(\tilde{x},\tilde{y})\sim\tilde{P}}{\mathbb{E}}[\tilde{y}(w\cdot\tilde{x})]$, which equals $w\cdot\mu$.
  • The Euclidean norm of the classifier vector $w$.
Lemma 9.
For any $w\in\mathbb{R}^{p}\setminus\{0\}$, there exists a positive constant $c$ such that
$\mathbb{P}_{(x,y)\sim P}\left[\mathrm{sign}(w\cdot x)\neq y\right]\leq\eta+2\exp\left(-c\frac{|w\cdot\mu|^{\alpha}}{\|w\|^{\alpha}}\right).$
Proof.
Following the proof of Chatterji and Long [6], we have
$\mathbb{P}_{(x,y)\sim P}\left[\mathrm{sign}(w\cdot x)\neq y\right]$ $=\mathbb{P}_{(x,y)\sim P}\left[y(w\cdot x)<0\right]$
$\leq\eta+\mathbb{P}_{(\tilde{x},\tilde{y})\sim\tilde{P}}\left[\tilde{y}(w\cdot\tilde{x})<0\right]$
$=\eta+\mathbb{P}_{(\tilde{x},\tilde{y})\sim\tilde{P}}\left[\left(\frac{w}{\|w\|}\cdot\tilde{y}\tilde{x}\right)<0\right]$
$=\eta+\mathbb{P}_{(\tilde{x},\tilde{y})\sim\tilde{P}}\left[\left(\frac{w}{\|w\|}\cdot\tilde{\xi}\right)<-\frac{w}{\|w\|}\cdot\mu\right]$
Applying Proposition 16, there exists a positive constant $c$ such that
$\mathbb{P}_{(\tilde{x},\tilde{y})\sim\tilde{P}}\left[\left(\frac{w}{\|w\|}\cdot\tilde{\xi}\right)<-\frac{w}{\|w\|}\cdot\mu\right]$ $\leq 2\exp\left(-c\frac{|w\cdot\mu|^{\alpha}}{\|w\|^{\alpha}}\right),$
which completes our proof. ∎
The next lemma provides concentration arguments for $z_{k}$.
Lemma 10.
For any $\alpha\in(0,2]$ and $\kappa\in(0,1)$, there exists a constant $c_{1}\geq 1$ such that, for any $c^{\prime}$, for all large enough $C$, with probability at least $1-\delta$, the following hold:
  1. 1.
    For any $k\in[n]$,
    $\frac{p}{c_{1}}\leq\|z_{k}\|^{2}\leq c_{1}p.$
  2. 2.
    For any $i\neq j\in[n]$,
    $|z_{i}\cdot z_{j}|\leq c^{\prime}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right).$
  3. 3.
    For any $k\in\mathcal{C}$,
    $|\mu\cdot z_{k}-\|\mu\|^{2}|<\frac{\|\mu\|^{2}}{2}$
  4. 4.
    For any $k\in\mathcal{N}$,
    $|\mu\cdot z_{k}-(-\|\mu\|^{2})|<\frac{\|\mu\|^{2}}{2}$
  5. 5.
    The number of noisy samples satisfies $|\mathcal{N}|\leq(\eta+c^{\prime})n$.
  6. 6.
    The samples are linearly separable.
The proof of this lemma is in Appendix A.2.
From here on, we will assume that samples satisfy all the conditions of Lemma 10. Next, we will prove that the ratio of the losses between any pair of points is bounded. In this proof, we use Lemma 11.
Lemma 11.
For any $s_{1},s_{2}\in\mathbb{R}$,
$\frac{1+\exp(s_{2})}{1+\exp(s_{1})}\leq\max\left(2,2\frac{\exp(-s_{1})}{\exp(-s_{2})}\right).$
First, we prove that the ratio of losses is bounded when the loss function is the exponential loss.
Lemma 12.
There exists a positive constant $c_{2}$ such that, for all large enough $C$, and any learning rate $\beta$ which satisfies
$\beta\leq\frac{1}{2}\left(c_{1}p+2nc_{1}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right)^{-1},$
for all iterations $t\geq 0$,
$\max_{i,j\in[n]}\left\{\frac{\exp(-\theta^{(t)}\cdot z_{i})}{\exp(-\theta^{(t)}\cdot z_{j})}\right\}\leq c_{2},$
where $c_{1}$ is a constant which satisfies Lemma 10.
Proof.
For simplicity, let $A_{t}$ be the ratio between exponential losses of the first and second samples for $t$ iterations:
$A_{t}=\frac{\exp(-\theta^{(t)}\cdot z_{1})}{\exp(-\theta^{(t)}\cdot z_{2})}.$
We will show that $A_{t}\leq 4c_{1}^{2}$ by using induction. When $t=0$, $A_{0}=1\leq 4c_{1}^{2}$. Thus, the base step holds. Assume that the inductive hypothesis holds for some iteration $t$. We shall now show that it must hold at iteration $t+1$.
$A_{t+1}$ $=\frac{\exp(-\theta^{(t+1)}\cdot z_{1})}{\exp(-\theta^{(t+1)}\cdot z_{2})}$
$=\frac{\exp\left(-(\theta^{(t)}-\beta\nabla R(\theta^{(t)}))\cdot z_{1}\right)}{\exp\left(-(\theta^{(t)}-\beta\nabla R(\theta^{(t)}))\cdot z_{2}\right)}$
$=A_{t}\frac{\exp\left(\beta\nabla R(\theta^{(t)})\cdot z_{1}\right)}{\exp\left(\beta\nabla R(\theta^{(t)})\cdot z_{2}\right)}$
$=A_{t}\frac{\exp\left(-\beta\sum_{k\in[n]}\frac{z_{k}\cdot z_{1}}{1+\exp(\theta^{(t)}\cdot z_{k})}\right)}{\exp\left(-\beta\sum_{k\in[n]}\frac{z_{k}\cdot z_{2}}{1+\exp(\theta^{(t)}\cdot z_{k})}\right)}$
$=A_{t}\frac{\exp\left(-\beta\frac{\|z_{1}\|^{2}}{1+\exp(\theta^{(t)}\cdot z_{1})}\right)}{\exp\left(-\beta\frac{\|z_{2}\|^{2}}{1+\exp(\theta^{(t)}\cdot z_{2})}\right)}\frac{\exp\left(-\beta\sum_{k\neq 1}\frac{z_{k}\cdot z_{1}}{1+\exp(\theta^{(t)}\cdot z_{k})}\right)}{\exp\left(-\beta\sum_{k\neq 2}\frac{z_{k}\cdot z_{2}}{1+\exp(\theta^{(t)}\cdot z_{k})}\right)}$
$=A_{t}\exp\left(-\beta\left(\frac{\|z_{1}\|^{2}}{1+\exp(\theta^{(t)}\cdot z_{1})}-\frac{\|z_{2}\|^{2}}{1+\exp(\theta^{(t)}\cdot z_{2})}\right)\right)$
×$\quad\times\exp\left(-\beta\left(\sum_{k\neq 1}\frac{z_{k}\cdot z_{1}}{1+\exp(\theta^{(t)}\cdot z_{k})}-\sum_{k\neq 2}\frac{z_{k}\cdot z_{2}}{1+\exp(\theta^{(t)}\cdot z_{k})}\right)\right)$
By Lemma 10, for any $k,i\neq j\in[n]$, there exists a constant $c_{1}$ such that
$\frac{p}{c_{1}}\leq\|z_{k}\|^{2}\leq c_{1}p,$
$|z_{i}\cdot z_{j}|\leq c_{1}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right).$
Thus,
$A_{t+1}$ $\leq A_{t}\exp\left(-\beta\left(\frac{p/c_{1}}{1+\exp(\theta^{(t)}\cdot z_{1})}-\frac{c_{1}p}{1+\exp(\theta^{(t)}\cdot z_{2})}\right)\right)$
×$\quad\times\exp\left(2\beta\sum_{k\in[n]}\frac{c_{1}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)}{1+\exp(\theta^{(t)}\cdot z_{k})}\right)$
$=A_{t}\exp\left(-\frac{\beta p}{c_{1}(1+\exp(\theta^{(t)}\cdot z_{2}))}\left(\frac{1+\exp(\theta^{(t)}\cdot z_{2})}{1+\exp(\theta^{(t)}\cdot z_{1})}-c_{1}^{2}\right)\right)$
×$\quad\times\exp\left(2\beta\sum_{k\in[n]}\frac{c_{1}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)}{1+\exp(\theta^{(t)}\cdot z_{k})}\right)$
Now we consider two disjoint cases.
Case 1 ($A_{t}<2c_{1}^{2}$):
$A_{t+1}$ $\leq A_{t}\exp\left(\frac{\beta c_{1}p}{1+\exp(\theta^{(t)}\cdot z_{2})}\right)\times\exp\left(2\beta nc_{1}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right)$
$\leq A_{t}\exp\left(\beta\left(c_{1}p+2nc_{1}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right)\right)$
Taking $\beta$ small enough that
$\beta\leq\frac{1}{2}\left(c_{1}p+2nc_{1}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right)^{-1},$
we have
$A_{t+1}\leq A_{t}\exp\left(\frac{1}{2}\right)\leq 2c_{1}^{2}\exp\left(\frac{1}{2}\right)<4c_{1}^{2}.$
Case 2 ($A_{t}\geq 2c_{1}^{2}$):
$A_{t+1}$ $=A_{t}\exp\left(-\frac{\beta p}{c_{1}(1+\exp(\theta^{(t)}\cdot z_{2}))}\left(\frac{1+\exp(\theta^{(t)}\cdot z_{2})}{1+\exp(\theta^{(t)}\cdot z_{1})}-c_{1}^{2}\right)\right)$
×$\quad\times\exp\left(2\beta c_{1}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\frac{1}{1+\exp(\theta^{(t)}\cdot z_{2})}\sum_{k\in[n]}\frac{1+\exp(\theta^{(t)}\cdot z_{2})}{1+\exp(\theta^{(t)}\cdot z_{k})}\right).$
By Lemma 11 and the induction hypothesis,
$A_{t+1}$ $\leq A_{t}\exp\left(-\frac{\beta c_{1}p}{1+\exp(\theta^{(t)}\cdot z_{2})}\right)$
×$\quad\times\exp\left(2\beta c_{1}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\frac{1}{1+\exp(\theta^{(t)}\cdot z_{2})}\sum_{k\in[n]}\max\left(2,2A_{t}\right)\right)$
$\leq A_{t}\exp\left(-\frac{\beta c_{1}p}{1+\exp(\theta^{(t)}\cdot z_{2})}\right)$
×$\quad\times\exp\left(2\beta c_{1}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\frac{n\max\left(2,8c_{1}^{2}\right)}{1+\exp(\theta^{(t)}\cdot z_{2})}\right)$
$\leq A_{t}\exp\left(-\frac{\beta c_{1}}{1+\exp(\theta^{(t)}\cdot z_{2})}\left(p-8c_{1}^{2}n\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right)\right)$
By Assumption (A3) and (A4), for large enough $C$,
$p-8c_{1}^{2}n\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)>0.$
Thus,
$A_{t+1}<A_{t}\leq 4c_{1}^{2}.$
This completes the proof of the inductive step. ∎
Next, we prove the bound on the ratio of losses when the loss function is the sigmoid loss.
Lemma 13.
There exists a positive constant $c_{3}$ such that, for all large enough $C$, and any learning rate $\beta$ which satisfies
$\beta\leq\frac{1}{2}\left(c_{1}p+2nc_{1}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right)^{-1},$
for all iterations $t\geq 0$,
$\max_{i,j\in[n]}\left\{\frac{1+\exp(\theta^{(t)}\cdot z_{j})}{1+\exp(\theta^{(t)}\cdot z_{i})}\right\}\leq c_{3},$
where $c_{1}$ is a constant which satisfies Lemma 10.
Proof.
This lemma is derived from Lemma 11 and Lemma 12. ∎
Now, we prove the lower bound of $\mu\cdot w$.
Lemma 14.
For any $\kappa\in(0,1)$, there exists a positive constant $c_{4}$ such that, for any large enough $C$, with probability at least $1-\delta$, the maximum margin weight vector $w$ satisfies,
$\mu\cdot w\geq\frac{\|w\|\|\mu\|^{2}}{c_{4}\sqrt{p}}.$
In the proof of Lemma 14, we use Lemma 15.
Lemma 15 (Soudry et al., 2018[27]).
For any linearly separable $\mathcal{S}$ and for all small enough learning rate $\beta$, we have
$\frac{w}{\|w\|}=\lim_{t\rightarrow\infty}\frac{\theta^{(t)}}{\|\theta^{(t)}\|}.$
Proof of Lemma 14.
$\mu\cdot\theta^{(t+1)}$ $=\mu\cdot\theta^{(t)}+\beta\sum_{k\in[n]}\frac{\mu\cdot z_{k}}{1+\exp(\theta^{(t)}\cdot z_{k})}$
$=\mu\cdot\theta^{(t)}+\beta\sum_{k\in\mathcal{C}}\frac{\mu\cdot z_{k}}{1+\exp(\theta^{(t)}\cdot z_{k})}+\beta\sum_{k\in\mathcal{N}}\frac{\mu\cdot z_{k}}{1+\exp(\theta^{(t)}\cdot z_{k})}.$
By Lemma 10,
$\mu\cdot\theta^{(t+1)}$ $\geq\mu\cdot\theta^{(t)}+\frac{\beta\|\mu\|^{2}}{2}\sum_{k\in\mathcal{C}}\frac{1}{1+\exp(\theta^{(t)}\cdot z_{k})}-\frac{3\beta\|\mu\|^{2}}{2}\sum_{k\in\mathcal{N}}\frac{1}{1+\exp(\theta^{(t)}\cdot z_{k})}$
$\geq\mu\cdot\theta^{(t)}+\frac{\beta\|\mu\|^{2}}{2}\sum_{k\in[n]}\frac{1}{1+\exp(\theta^{(t)}\cdot z_{k})}-2\beta\|\mu\|^{2}\sum_{k\in\mathcal{N}}\frac{1}{1+\exp(\theta^{(t)}\cdot z_{k})}.$
By $|\mathcal{N}|\leq(\eta+c^{\prime})n$ and Lemma 13,
$\sum_{k\in\mathcal{N}}\frac{1}{1+\exp(\theta^{(t)}\cdot z_{k})}$ $\leq c_{3}(\eta+c^{\prime})n\min_{k\in[n]}\frac{1}{1+\exp(\theta^{(t)}\cdot z_{k})}$
$\leq c_{3}(\eta+c^{\prime})\sum_{k\in[n]}\frac{1}{1+\exp(\theta^{(t)}\cdot z_{k})},$
where $c_{3}$ is the constant from Lemma 13. Recalling that $\eta\leq\frac{1}{C}$ and $c^{\prime}$ is an arbitrary positive constant, for large enough $C$ and small enough $c^{\prime}$,
$\sum_{k\in\mathcal{N}}\frac{1}{1+\exp(\theta^{(t)}\cdot z_{k})}\leq\frac{1}{8}\sum_{k\in[n]}\frac{1}{1+\exp(\theta^{(t)}\cdot z_{k})}.$
Thus, we have
$\mu\cdot\theta^{(t+1)}$ $\geq\mu\cdot\theta^{(t)}+\frac{\beta\|\mu\|^{2}}{4}\sum_{k\in[n]}\frac{1}{1+\exp(\theta^{(t)}\cdot z_{k})}.$
By using this inequality repeatedly and $\theta^{(0)}=0$,
$\mu\cdot\theta^{(t+1)}$ $\geq\frac{\beta\|\mu\|^{2}}{4}\sum_{m=0}^{t}\sum_{k\in[n]}\frac{1}{1+\exp(\theta^{(m)}\cdot z_{k})},$
$\|w\|\frac{\mu\cdot\theta^{(t+1)}}{\|\theta^{(t+1)}\|}$ $\geq\|w\|\frac{\beta\|\mu\|^{2}\sum_{m=0}^{t}\sum_{k\in[n]}\frac{1}{1+\exp(\theta^{(m)}\cdot z_{k})}}{4\|\theta^{(t+1)}\|}.$
By taking the large-$t$ limit and using Lemma 15,
$\mu\cdot w\geq\beta\|w\|\|\mu\|^{2}\lim_{t\rightarrow\infty}\frac{\sum_{m=0}^{t}\sum_{k\in[n]}\frac{1}{1+\exp(\theta^{(m)}\cdot z_{k})}}{4\|\theta^{(t+1)}\|}$ (1)
By definition of gradient descent iterations,
$\|\theta^{(t+1)}\|$ $=\left\|\sum_{m=0}^{t}\beta\nabla R(\theta^{(m)})\right\|$
$\leq\beta\sum_{m=0}^{t}\|\nabla R(\theta^{(m)})\|$
$\leq\beta\sum_{m=0}^{t}\left\|\sum_{k\in[n]}\frac{-z_{k}}{1+\exp(\theta^{(m)}\cdot z_{k})}\right\|$
$\leq\beta c_{1}\sqrt{p}\sum_{m=0}^{t}\sum_{k\in[n]}\frac{1}{1+\exp(\theta^{(m)}\cdot z_{k})}.$
With inequality (1), we have
$\mu\cdot w\geq\frac{\|w\|\|\mu\|^{2}}{4c_{1}\sqrt{p}},$
which completes our proof. ∎
By Lemma 9 and Lemma 14, we have Theorem 4.
5 Conclusion
Our research extends the analysis of benign overfitting in binary classification problems to heavy-tailed input distributions, specifically $\alpha$ sub-exponential distributions where $\alpha\in(0,2]$. The main findings of this study are:
  1. 1.
    We derived generalization bounds for maximum margin classifiers in this heavy-tailed setting, showing that benign overfitting can occur under certain conditions on dimensionality $p$ and the feature vector magnitude $\|\mu\|$.
  2. 2.
    Our results demonstrate that as the number of dimensions $p$ increases and the feature vector magnitude $\|\mu\|$ scales as $\Theta(p^{d})$ for $d\in(1/4,1/2]$, the misclassification error approaches the noise level $\eta$.
  3. 3.
    In the context of the noisy rare-weak model, our upper bounds suggest that the maximum margin classifier can succeed arbitrarily close to the known impossibility threshold of $s=O(\sqrt{p})$.
These findings significantly contribute to our understanding of benign overfitting by showing that the phenomenon is not limited to sub-gaussian distributions but extends to heavier-tailed inputs as well. This research bridges a gap between theory and practice, as real-world data often exhibit heavier tails than the gaussian distribution.
Our work opens up several avenues for future research:
  1. 1.
    Investigation of benign overfitting in even heavier-tailed distributions, such as those with polynomial tails.
  2. 2.
    Extension of the analysis to multi-class classification problems with heavy-tailed inputs.
  3. 3.
    Exploration of the implications of these findings for deep learning models, which often deal with high-dimensional, heavy-tailed data.
  4. 4.
    Development of new learning algorithms that explicitly leverage the properties of heavy-tailed distributions to achieve better generalization in high-dimensional settings.
In conclusion, this study provides a significant step towards understanding the phenomenon of benign overfitting in more realistic data settings. By extending the theory to heavy-tailed distributions, we have broadened the applicability of benign overfitting results to a wider range of practical scenarios, potentially impacting the design and analysis of machine learning algorithms for complex, real-world data.
Acknowledgements
We would like to thank our colleagues for their valuable feedback and suggestions that helped improve this work. This research was supported in part by RIKEN AIP and JSPS KAKENHI (JP22K03439).
Appendix A Proof of lemmas
A.1 Concentration inequality
In this section, we introduce the concentration inequalities for $\alpha$ sub-exponential random variables. In our proof, we apply the following two concentration inequalities.
Proposition 16 (A special case of Theorem 1.5 in [12]).
Let $\alpha\in(0,2]$ and $K$ be a positive constant and $a\in\mathbb{R}^{n}$ be a constant vector. Let $X_{1},\ldots,X_{n}$ be independent mean-zero random variables satisfying $\|X_{i}\|_{\psi_{\alpha}}\leq K$. Then, there exists a positive constant $c_{\alpha}$ such that for any $t>0$ it holds
$\mathbb{P}\left[\left|\sum_{i=1}^{n}a_{i}X_{i}\right|\geq t\right]\leq 2\exp\left(-\frac{1}{c_{\alpha}}\frac{t^{\alpha}}{K^{\alpha}\|a\|^{\alpha}}\right).$
This is a special case of Theorem 1.5 of Götze et al. [12].
Theorem 17 (Extended Hanson-Wright inequality [26]).
Let $\alpha\in(0,2]$ and $K$ be a positive constant. Let $X_{1},\ldots,X_{n}$ be independent mean-zero random variables such that $\|X_{i}\|_{\psi_{\alpha}}\leq K$, the corresponding random vector $X$ be $(X_{1},\ldots,X_{n})^{T}$ and $A\in\mathbb{R}^{n\times n}$ be a symmetric matrix. Then, there exists a positive constant $c_{\alpha}$ such that it holds
$\mathbb{P}[|X^{T}AX-\mathbb{E}[X^{T}AX]|\geq t]\leq 2\exp\left(-\frac{1}{c_{\alpha}}\min\left(\frac{t^{2}}{K^{4}\|A\|_{\mathrm{HS}}^{2}},\left(\frac{t}{K^{2}\|A\|_{\mathrm{op}}}\right)^{\frac{\alpha}{2}}\right)\right)$
for any $t\geq 0$.
A.2 Proof of Lemma 10
In this section, we prove Lemma 10 by using concentration inequalities from Section A.1. We assume Assumptions (A1)-(A4) hold, and decompose Lemma 10 into six different parts. We prove that each separate lemma holds with probability at least $61-\delta/$ 6.
Lemma 18.
For any $\alpha\in(0,2]$ and $\kappa\in(0,1)$, there exists a constant $c\geq 1$ such that, for all large enough $C$, with probability at least $61-\delta/$ 6, for any $k\in[n]$,
$\frac{p}{c}\leq\|z_{k}\|^{2}\leq cp.$
Proof.
By Theorem 17 with $A=I$, there exists a positive constant $c$ such that
$\mathbb{P}\left[\left|\|\xi_{k}\|^{2}-\mathbb{E}[\|\xi_{k}\|^{2}]\right|\geq t\right]\leq 2\exp\left(-\frac{1}{c}\min\left(\frac{t^{2}}{p},t^{\frac{\alpha}{2}}\right)\right).$
By setting $t=\frac{\kappa p}{2}$
$2\exp\left(-\frac{1}{c}\min\left(\frac{t^{2}}{p},t^{\frac{\alpha}{2}}\right)\right)=2\exp\left(-\frac{1}{c}\min\left(\left(\frac{\kappa}{2}\right)^{2}p,\left(\frac{\kappa}{2}\right)^{\frac{\alpha}{2}}p^{\frac{\alpha}{2}}\right)\right)$
By Assumption (A3), we have $p\geq C\left(\log\frac{n}{\delta}\right)^{\frac{2}{\alpha}}$. There exists a large enough constant $C$ such that
$2\exp\left(-\frac{1}{c}\min\left(\left(\frac{\kappa}{2}\right)^{2}p,\left(\frac{\kappa}{2}\right)^{\frac{\alpha}{2}}p^{\frac{\alpha}{2}}\right)\right)\leq\frac{\delta}{6n}.$
Thus,
$\mathbb{P}\left[\left|\|\xi_{k}\|^{2}-\mathbb{E}[\|\xi_{k}\|^{2}]\right|\geq\frac{\kappa p}{2}\right]\leq\frac{\delta}{6n}.$ (2)
Recalling the assumption $\mathbb{E}[\|q\|^{2}]\geq\kappa p$, we have
$\mathbb{E}[\|\xi_{k}\|^{2}]=\mathbb{E}\left[\|z_{k}-\mathbb{E}[z]\|^{2}\right]=\mathbb{E}[\|q\|^{2}]\geq\kappa p.$
Let $\{\xi_{kj}\}_{j=1}^{p}$ be elements of $\xi_{k}$. By $\|\xi_{kj}\|_{\psi_{\alpha}}\leq 1$ for each $j$,
[…]
$=\frac{4}{\alpha}\Gamma\left(\frac{2}{\alpha}\right).$
Thus, $\mathbb{E}[\|\xi_{k}\|^{2}]\leq\frac{4p}{\alpha}\Gamma\left(\frac{2}{\alpha}\right)$. Because of this and (2), with probability at least $1-\frac{\delta}{6n}$,
$\frac{\kappa p}{2}\leq\|\xi_{k}\|^{2}\leq\left(\frac{4}{\alpha}\Gamma\left(\frac{2}{\alpha}\right)+\frac{\kappa}{2}\right)p.$
Suppose $k\in\mathcal{C}$, and let $\xi_{k}=z_{k}-\mu$. Then, the following inequalities hold.
  1. 1.
    $\|z_{k}-\mu\|^{2}\leq 2\|z_{k}\|^{2}+2\|\mu\|^{2}$.
  2. 2.
    $\|\mu\|^{2}<\frac{p}{C}$ by Assumption (A3).
  3. 3.
    $\|\xi_{k}\|^{2}=\|z_{k}-\mu\|^{2}\geq\frac{\kappa p}{2}$ with probability at least $1-\frac{\delta}{6}$.
Combining these inequalities, we obtain:
$\|z_{k}\|^{2}$ $\geq\frac{1}{2}\|z_{k}-\mu\|^{2}-\|\mu\|^{2}$
$\geq\frac{1}{2}\left(\frac{\kappa p}{2}\right)-\left(\frac{p}{C}\right)$
$=\frac{\kappa p}{4}-\frac{p}{C}$
For sufficiently large $C$, we can ensure that $\frac{p}{C}<\frac{\kappa p}{8}$. Thus,
$\|z_{k}\|^{2}>\frac{\kappa p}{4}-\frac{\kappa p}{8}=\frac{\kappa p}{8}$ (3)
Therefore, with probability at least $1-\frac{\delta}{6n}$, we have $\|z_{k}\|^{2}>\frac{\kappa p}{8}$ for sufficiently large $C$. Again by $\|z_{k}\|^{2}\leq 2\|z_{k}-\mu\|^{2}+2\|\mu\|^{2}$,
$\|z_{k}\|^{2}$ $\leq 2\|z_{k}-\mu\|^{2}+2\|\mu\|^{2}$
$\leq 2\left(\frac{4}{\alpha}\Gamma\left(\frac{2}{\alpha}\right)+\frac{\kappa}{2}\right)p+2\|\mu\|^{2}$
$<2\left(\frac{4}{\alpha}\Gamma\left(\frac{2}{\alpha}\right)+\frac{\kappa}{2}\right)p+\frac{2p}{C}$
$<\left(\frac{8}{\alpha}\Gamma\left(\frac{2}{\alpha}\right)+\kappa+2\right)p$
Therefore, setting $c=\max\left(\frac{8}{\kappa},\frac{8}{\alpha}\Gamma\left(\frac{2}{\alpha}\right)+\kappa+2\right)$, we have
$\frac{p}{c}\leq\|z_{k}\|^{2}\leq cp$
for any $k\in\mathcal{C}$. A similar argument holds for $k\in\mathcal{N}$. ∎
Lemma 19.
There exists $c\geq 1$ such that, for any large enough $C$, with probability at least $1-\frac{\delta}{6}$, for any $i\neq j\in[n]$,
$|z_{i}\cdot z_{j}|\leq c\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right).$
Proof.
Applying Theorem 17 as in the proof of Lemma 18 and using the union bound method, we have
$\mathbb{P}[\exists i\in[n],\|\xi_{i}\|\geq\sqrt{p}]\leq\frac{\delta}{24}.$
For any pair $i,j\in[n]$, we have
$\mathbb{P}[|\xi_{i}\cdot\xi_{j}|\geq t]\leq\mathbb{P}[|\xi_{i}\cdot\xi_{j}|\geq t\mid\|\xi_{j}\|\leq\sqrt{p}]+\mathbb{P}[\|\xi_{j}\|>\sqrt{p}].$ (4)
Regarding $\xi_{j}$ as fixed, by Proposition 16 there exists a positive constant $c_{2}$ such that
$\mathbb{P}[|\xi_{i}\cdot\xi_{j}|\geq t]=\mathbb{P}\left[\frac{\xi_{j}}{\|\xi_{j}\|}\cdot\xi_{i}\geq\frac{t}{\|\xi_{j}\|}\right]\leq 2\exp\left(-c_{2}\frac{t^{\alpha}}{\|\xi_{j}\|^{\alpha}}\right).$
Therefore,
$\mathbb{P}[|\xi_{i}\cdot\xi_{j}|\geq t\mid\|\xi_{j}\|\leq\sqrt{p}]$ $\leq 2\exp\left(-c_{2}\frac{t^{\alpha}}{p^{\frac{\alpha}{2}}}\right)$
$\mathbb{P}[|\xi_{i}\cdot\xi_{j}|\geq t]$ $\leq 2\exp\left(-c_{2}\frac{t^{\alpha}}{p^{\frac{\alpha}{2}}}\right)+\mathbb{P}[\|\xi_{j}\|>\sqrt{p}].$
By the union bound method,
$\mathbb{P}\left[\exists i\neq j\in[n],|\xi_{i}\cdot\xi_{j}|\geq t\right]\leq 2n^{2}\exp\left(-c_{2}\frac{t^{\alpha}}{p^{\frac{\alpha}{2}}}\right)+\mathbb{P}[\exists j\in[n],\|\xi_{j}\|>\sqrt{p}]$
Setting $t=c_{3}\left(p^{\frac{\alpha}{2}}\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}$, for a large enough $c_{3}$, we have
$\mathbb{P}\left[\exists i\neq j\in[n],|\xi_{i}\cdot\xi_{j}|\geq c_{3}\left(p^{\frac{\alpha}{2}}\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right]\leq\frac{\delta}{24}+\mathbb{P}[\exists j\in[n],\|\xi_{j}\|>\sqrt{p}]$
Together with the inequality (4), we have,
$\mathbb{P}\left[\exists i\neq j\in[n],|\xi_{i}\cdot\xi_{j}|\geq c_{3}\left(p^{\frac{\alpha}{2}}\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right]\leq\frac{\delta}{12}.$ (5)
By Proposition 16, there exists a constant $c_{4}$ such that
$\mathbb{P}[|\mu\cdot z_{k}|>\|\mu\|^{2}]$ $=\mathbb{P}\left[\left|\frac{\mu}{\|\mu\|}\cdot z_{k}\right|>\|\mu\|\right]\leq 2\exp\left(-c_{4}\|\mu\|^{\alpha}\right).$
By Assumption (A4), for large enough $C$ we have
$\mathbb{P}[|\mu\cdot z_{k}|>\|\mu\|^{2}]\leq\frac{\delta}{12n}.$
By taking a union bound, we have
$\mathbb{P}[\exists k,|\mu\cdot z_{k}|>\|\mu\|^{2}]\leq\frac{\delta}{12}.$ (6)
Due to inequalities (5) and (6), with probability at least $1-\frac{\delta}{6}$,
$|z_{i}\cdot z_{j}|$ $=\Big{|}(z_{i}-\mathbb{E}[z_{i}])\cdot(z_{j}-\mathbb{E}[z_{j}])-\mathbb{E}[z_{i}]\cdot\mathbb{E}[z_{j}]+\mathbb{E}[z_{i}]\cdot z_{j}+\mathbb{E}[z_{j}]\cdot z_{i}\Big{|}$
$=\left|\xi_{i}\cdot\xi_{j}-\|\mu\|^{2}+\mu\cdot z_{j}+\mu\cdot z_{i}\right|$
$\leq|\xi_{i}\cdot\xi_{j}|+\|\mu\|^{2}+|\mu\cdot z_{j}|+|\mu\cdot z_{i}|$
$\leq 3\|\mu\|^{2}+c\left(p^{\frac{\alpha}{2}}\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}$
Lemma 20.
For any large enough $C$, with probability at least $1-\frac{\delta}{6}$, for $k\in\mathcal{C}$,
$|\mu\cdot z_{k}-\|\mu\|^{2}|\leq\frac{\|\mu\|^{2}}{2}.$
Lemma 21.
For any large enough $C$, with probability at least $1-\frac{\delta}{6}$, for $k\in\mathcal{N}$,
$|\mu\cdot z_{k}-(-\|\mu\|^{2})|\leq\frac{\|\mu\|^{2}}{2}.$
Lemma 20 and Lemma 21 can be proven by the same logic. We will prove Lemma 20.
Proof.
If $k\in\mathcal{C}$,
$\mu\cdot z_{k}-\|\mu\|^{2}=\mu\cdot\xi_{k}.$
Since the exponential Orlicz norm of $\xi_{k}$ is at most 1, by Proposition 16, there exists a positive constant $c$ such that
$\mathbb{P}\left[\left|\mu\cdot z_{k}-\|\mu\|^{2}\right|\geq\frac{\|\mu\|^{2}}{2}\right]$ $=\mathbb{P}\left[\left|\frac{\mu}{\|\mu\|}\cdot\xi_{k}\right|\geq\frac{\|\mu\|}{2}\right]$
$\leq 2\exp\left(-c\frac{\|\mu\|^{\alpha}}{2^{\alpha}}\right).$
By Assumption (A4), for large enough constant $C$ we have
$\mathbb{P}\left[\left|\mu\cdot z_{k}-\|\mu\|^{2}\right|\geq\frac{\|\mu\|^{2}}{2}\right]\leq\frac{\delta}{6n}.$
Taking a union bound, we have
$\mathbb{P}\left[\exists k\in\mathcal{C},\left|\mu\cdot z_{k}-\|\mu\|^{2}\right|\geq\frac{\|\mu\|^{2}}{2}\right]\leq\frac{\delta}{6},$
which completes our proof. ∎
Lemma 22.
For any $c^{\prime}>0$, for any large enough $C$, with probability at least $1-\frac{\delta}{6}$, the number of noisy samples satisfies $|\mathcal{N}|\leq(\eta+c^{\prime})n$.
Proof.
$\mathbb{E}\left[|\mathcal{N}|\right]=\sum_{k=1}^{n}\mathbb{E}\left[1_{\{y_{k}\neq\tilde{y}_{k}\}}\right]=\sum_{k=1}^{n}\mathbb{P}[y_{k}\neq\tilde{y}_{k}]=n\eta.$
By Hoeffding’s inequality,
$\mathbb{P}\left[|\mathcal{N}|\geq(\eta+c^{\prime})n\right]$ $=\mathbb{P}\left[\frac{1}{n}\sum_{k=1}^{n}\left(1_{\{y_{k}\neq\tilde{y}_{k}\}}-\mathbb{E}[1_{\{y_{k}\neq\tilde{y}_{k}\}}]\right)\geq c^{\prime}\right]$
$\leq 2\exp\left(-2{c^{\prime}}^{2}n\right)$
$\leq\frac{\delta}{6}.$
The last inequality holds due to Assumption (A2). ∎
Lemma 23.
If the following conditions hold, for any large enough $C$, $\{(x_{k},y_{k})\}_{k=1}^{n}$ are linearly separable.
  1. 1.
    There exists a positive constant $c$ such that for any $k\in[n]$
    $\frac{p}{c}\leq\|z_{k}\|^{2}\leq cp.$
  2. 2.
    There exists a positive constant $c$ such that for any $i\neq j\in[n]$
    $|z_{i}\cdot z_{j}|\leq c\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right).$
Proof.
Let $v$ be $\sum_{i\in[n]}z_{i}$. For each $k\in[n]$ and any $\delta>0$,
$y_{k}v\cdot x_{k}$ $=\sum_{i\in[n]}z_{i}\cdot z_{k}$
$=\|z_{k}\|^{2}+\sum_{i\neq k}z_{i}\cdot z_{k}$
$\geq\|z_{k}\|^{2}-\sum_{i\neq k}|z_{i}\cdot z_{k}|$
$\geq\frac{p}{c}-cn\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)$
$\geq\frac{p}{c}-2cn\max\left(\|\mu\|^{2},\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)$
$=\frac{1}{c}\left(p-2c^{2}n\max\left(\|\mu\|^{2},\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right)$
By Assumption (A3) and (A4), for large enough $C$ we have
$y_{k}v\cdot x_{k}>0,$
which completes our proof. ∎