### v2

# Benign Overfitting under Learning Rate Conditions for

Abstract
This paper investigates the phenomenon of benign overfitting in binary classification problems with heavy-tailed input distributions, extending the analysis of maximum margin classifiers to $\alpha$ sub-exponential distributions ($\alpha\in(0,2]$). This generalizes previous work focused on sub-gaussian inputs. We provide generalization error bounds for linear classifiers trained using gradient descent on unregularized logistic loss in this heavy-tailed setting. Our results show that, under certain conditions on the dimensionality $p$ and the distance between the centers of the distributions, the misclassification error of the maximum margin classifier asymptotically approaches the noise level, the theoretical optimal value. Moreover, we derive an upper bound on the learning rate $\beta$ for benign overfitting to occur and show that as the tail heaviness of the input distribution $\alpha$ increases, the upper bound on the learning rate decreases. These results demonstrate that benign overfitting persists even in settings with heavier-tailed inputs than previously studied, contributing to a deeper understanding of the phenomenon in more realistic data environments.
1 Introduction
In the field of machine learning, a phenomenon that contradicts the long-standing intuition of statistical learning theory has been garnering attention. This phenomenon is called benign overfitting. According to conventional theory, when a model excessively fits the training data, its generalization performance on unseen data was expected to decline. However, experiments using deep neural networks have revealed that models that perfectly fit noisy training data surprisingly demonstrate good performance on unseen data as well [37, 2].
This phenomenon suggests a significant gap between machine learning theory and practice, attracting the attention of researchers. To deepen our understanding of benign overfitting, studies have been conducted in simpler statistical settings that are more amenable to theoretical analysis, such as linear regression [14, 1, 26, 27, 31, 10, 9], sparse linear regression [16, 8, 18, 33], logistic regression [24, 7, 21, 25, 34, 23, 12, 38], and kernel-based estimators [3, 22, 19, 20]. These studies are rapidly advancing our understanding of the conditions and mechanisms under which benign overfitting occurs.
In the context of binary classification, a standard mixture model is often used to study benign overfitting [7, 12, 38]. This model involves classifying well-separated data with adversarially corrupted labels, assuming the input distribution is sub-gaussian. However, benign overfitting in settings with input distributions heavier than sub-gaussian, that is, settings more robust to input variations, has not been extensively discussed.
Figure 1: Boxplot of estimated tail index $\xi$ for feature vector components extracted from the intermediate layers of a CNN with ReLU activation, trained on various datasets (CIFAR-10 [17], CIFAR-100 [17], Fashion-MNIST [36], SVHN [28]). The tail index $\xi$ represents the heaviness of the distribution tails, with smaller values indicating heavier tails. The Gaussian and Exponential distributions are included for comparison purposes and were not passed through the CNN. The results indicate that the feature vectors for certain datasets, have heavier-tailed distributions than the Gaussian distribution. Further details are found in Appendix C.1.
Figure 2: Training and test errors versus dimension $p$ for a maximum margin classifier. $n_{\mathrm{train}}=200$, $n_{\mathrm{test}}=1000$, $p$ ranges from 100 to 1500. Data is generated from a heavy-tailed setting using a generalized normal distribution, as detailed in Section 2.3 and Appendix C.2. The shape parameters are $\gamma=0.25,0.5,2$, with variance normalized to 1. Noise level $\eta$ is $0.05$ (dotted line). Solid and dashed lines show training and test errors, respectively, with $95$% confidence intervals as error bars over 50 trials. Training error remains near zero, while test error stabilizes around the noise level as $p$ increases.
Our numerical experiments indicate that feature vectors in convolutional neural networks (CNNs) with ReLU activation often exhibit distributions with tails heavier than sub-gaussian. The tail index $\xi$ can be intuitively understood as a parameter that characterizes the heaviness of the tails of a distribution. Specifically, for large values of $t$, the tail probability can be approximated as $\mathbb{P}[|X|>t]\simeq a\cdot\exp(-b\cdot t^{\xi})$, where smaller values of $\xi$ correspond to heavier tails in the distribution. Figure 1 indicates that these feature vectors have significantly heavy-tailed components. This finding emphasizes the necessity to extend benign overfitting analysis to accommodate a wider range of distributional settings. A more detailed explanation is provided in Appendix C.1.
Moreover, our numerical experiments suggest that benign overfitting can occur even in mixture model settings where the distributions have heavier tails than the normal distribution, as seen in Figure 2. Further details are available in Appendix C.2. This motivates the exploration of benign overfitting in more general distributional frameworks.
In this work, we focus on binary classification tasks where the input distribution is $\alpha$ sub-exponential with $\alpha\in(0,2]$, implying tails heavier than sub-gaussian. We aim to establish generalization error bounds that demonstrate benign overfitting for a linear classifier trained using gradient descent on the unregularized logistic loss. Moreover, we derive an upper bound on the learning rate, a factor previously unexamined in this context, which plays a crucial role in demonstrating benign overfitting.
In this paper, we focus on the setting of Chatterji and Long [7], a pioneering work of benign overfitting theory on a simple model. Their results are extended to the heavy-tailed setting, and a more detailed discussion on the learning rate is provided.
1.1 Related works
Benign overfitting in classification: Most related to our work is the theoretical analysis of benign overfitting in classification settings. This line of research aims to understand why classifiers that perfectly fit noisy training data can still generalize well to unseen data. Wang and Thrampoulidis [35] studied a setting where the two classes are symmetric mixture of Gaussian (or sub-gaussian) distributions, without label noise. Chatterji and Long [7] studied overparameterized linear logistic regression on sub-gaussian mixture models with label flipping noise. They showed how gradient descent can train these models to achieve nearly optimal population risk. Cao et al. [6] extended this work, tightening the upper bound in the case without label flipping noise and establishing a matching lower bound for overparameterized maximum margin interpolators. Wang et al. [34] extended the analysis of maximum margin classifiers to multiclass classification in overparameterized settings. Frei et al. [12] proved that a two-layer fully connected neural network exhibits benign overfitting under certain conditions, such as well-separated log-concave distribution and smooth activation function. Cao et al. [5] focused on the benign overfitting of two-layer convolutional neural networks.
2 Preliminaries
In this section, we introduce the definition of $\alpha$ sub-exponential random variables, the assumptions on the data generation process, and the maximum margin algorithm we consider.
2.1 Notation
In this paper, we use the notation $[n]$ to denote the set $\{1,2,\ldots,n\}$ for a positive integer $n$. For a vector $x$, we use $\|x\|$ to denote its $\ell^{2}$ norm. For a matrix $A$, we use $\|A\|_{\mathrm{HS}}$ to denote its Hilbert–Schmidt norm and $\|A\|_{\mathrm{op}}$ to denote its operator norm. We use $s_{k}(A)$ to denote the $k$-th largest singular value of $A$. We use $O(\cdot)$ and $\Theta(\cdot)$ to refer to big-O and big-Theta notation.
2.2 $\alpha$ sub-exponential random variable
$\alpha$ sub-exponential random variables are random variables which have exponential type tails.
Definition 1 ($\alpha$ sub-exponential random variable, [29]).
A random variable $X$ is called $\alpha$ sub-exponential if there is a positive constant $c_{\alpha}$ such that it holds
$\mathbb{P}\left[|X-\mathbb{E}[X]|\geq t\right]\leq 2\exp\left(-\frac{t^{\alpha}}{c_{\alpha}}\right)$
for any $t>0$. This is equivalent to having a finite exponential Orlicz norm:
$\|X\|_{\psi_{\alpha}}:=\inf\left\{t>0:\mathbb{E}\left[\exp\left(\frac{|X|^{\alpha}}{t^{\alpha}}\right)\right]\leq 2\right\}<\infty.$
If $\alpha=2$, we call the distribution sub-gaussian. If $\alpha=1$, we call it sub-exponential. Here is an example of an $\alpha$ sub-exponential distribution:
Example 2 (Generalized normal distribution).
The probability density function of the generalized normal distribution is defined as:
$f(x;x_{0},\sigma,\gamma)=\frac{\gamma}{2\sigma\Gamma(1/\gamma)}\exp\left(-\left|\frac{x-x_{0}}{\sigma}\right|^{\gamma}\right)$
where $\Gamma$ denotes the gamma function, $x_{0}$ is the location parameter, $\sigma>0$ is the scale parameter, and $\gamma>0$ is the shape parameter. Let $X$ be a random variable following the generalized normal distribution with location parameter $x_{0}=0$. Then $\gamma$ is the maximum value of $\alpha$ such that $\|X\|_{\psi_{\alpha}}$ is finite, and its exponential Orlicz norm is given by
$\|X\|_{\psi_{\gamma}}=\frac{\sigma}{(1-2^{-\gamma})^{1/\gamma}}.$
2.3 Data generation process
We consider a heavy-tailed setting for binary classification, which is a relaxed setting of a standard mixture model setting (Chatterji and Long, 2021 [7]; Frei et al., 2022 [12]). We first define a “clean” distribution $\tilde{P}$ and then define the target distribution $P$ based on $\tilde{P}$:
  1. 1.
    Sample a “clean” label $\tilde{y}\in\{\pm 1\}$ uniformly at random, $\tilde{y}\sim\mathrm{Uniform}(\{\pm 1\})$.
  2. 2.
    Sample $q\sim P_{\mathrm{clust}}$ that satisfies:
    • $P_{\mathrm{clust}}:=P_{\mathrm{clust}}^{(1)}\times\cdots\times P_{\mathrm{clust}}^{(p)}$ is an arbitrary product distribution on $\mathbb{R}^{p}$ whose marginals are all mean-zero with the exponential Orlicz norm at most $1$, i.e., $\|X\|_{\psi_{\alpha}}\leq 1$ if $X\sim P_{\rm clust}^{(j)}$.
    • For some $\kappa>0$, it holds that
      $\mathbb{E}_{q\sim P_{\mathrm{clust}}}[\|q\|^{2}]\geq\kappa p.$
  3. 3.
    For an arbitrary orthogonal matrix $U\in\mathbb{R}^{p\times p}$ and $\mu\in\mathbb{R}^{p}$, generate $\tilde{x}=Uq+\tilde{y}\mu$.
  4. 4.
    Let $\tilde{P}$ be the distribution of $(\tilde{x},\tilde{y})$.
  5. 5.
    For $\eta\in[0,1]$, let $P$ be an arbitrary distribution on $\mathbb{R}^{p}\times\{\pm 1\}$ that satisfies:
    • All marginal distributions of $P$ are the same as $\tilde{P}$.
    • Total variation between $P$ and $\tilde{P}$ is at most $\eta$.
Let $\mathcal{S}:=\{(x_{1},y_{1}),\cdots,(x_{n},y_{n})\}$ be samples drawn according to $P$.
The reason we assume the $\alpha$ sub-exponential norm of each component is at most $1$ is only for simplifying the proofs and does not affect the main results of the paper, since rescaling the data does not affect the accuracy of the maximum margin algorithm.
This setting is a modification of Chatterji and Long’s framework [7], where we have replaced the sub-gaussian norm with an exponential Orlicz norm. Moreover, it can be observed that when $\alpha=2$, this setting encompasses the original framework.
2.4 Maximum margin algorithm
We consider a linear classifier that takes the form $\mathrm{sign}(\theta\cdot x)$ trained by gradient descent as
$\theta^{(t+1)}=\theta^{(t)}-\beta\nabla R(\theta^{(t)})$
$\quad\mathrm{where}\quad R(\theta):=\sum_{i=1}^{n}\log(1+\exp(-y_{i}\theta\cdot x_{i})),$
where $\beta$ is the learning rate. In Soudry et al., 2018 [30], they prove that if the dataset is linearly separable, in the large-$t$ limit, the normalized parameter of this classifier converges to the hard margin predictor:
$\lim_{t\rightarrow\infty}\frac{\theta^{(t)}}{\|\theta^{(t)}\|}=\frac{w}{\|w\|},$
$w:=\underset{u\in\mathbb{R}^{p}}{\mathrm{argmin}}\|u\|$
$\text{such that }y_{i}(u\cdot x_{i})\geq 1,\text{ for any }i\in[n].$
They have proved this for a class of loss functions with certain smoothness.
2.5 Assumptions
We assume that $\alpha$ and $\kappa$ are fixed constants. Let $X=[y_{1}x_{1},\cdots,y_{n}x_{n}]$, where $\{(x_{k},y_{k})\}_{k=1}^{n}$ are samples drawn according from the distribution $P$. We will prove the main theorem and corollaries under the following assumptions with a sufficiently large constant $C$ depending only on $\alpha$ and $\kappa$.
  1. (A1)
    The failure probability satisfies $\delta\in(0,\frac{1}{C})$,
  2. (A2)
    The number of samples satisfies $n\geq C\log\frac{1}{\delta}$,
  3. (A3)
    The dimension satisfies
    $p\geq C\max\left(\|\mu\|^{2}n,n^{2}\left(\log\frac{n}{\delta}\right)^{\frac{2}{\alpha}}\right),$
  4. (A4)
    The norm of the mean satisfies $\|\mu\|\geq C\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}$,
  5. (A5)
    The learning rate satisfies
    $\beta\leq$ $\min\Bigg{(}8(s_{1}(X))^{-2},$
    $\frac{1}{c_{2}}\left(p+2n\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right)^{-1}\Bigg{)},$
    where $c_{2}=2\max\left(\frac{8}{\kappa},\frac{8}{\alpha}\Gamma\left(\frac{2}{\alpha}\right)+\kappa+2\right).$
When $\alpha=2$, assumptions (A1)-(A4) correspond to those in Chatterji and Long [7].
Due to assumption (A1), if we require a lower failure probability, $C$ must be set large, which in turn requires tighter lower bounds for $n$, $p$ and $\|\mu\|$ in assumptions (A2)-(A4). Moreover, as the tail heaviness of the distribution grows (i.e., as $\alpha$ decreases), the lower bounds for $n$, $p$, and $\|\mu\|$ become tighter, and the upper bound on the learning rate $\beta$ becomes more restrictive.
Example 3 (Generalized noisy rare-weak model).
For any $\alpha\in(0,2]$, the model described above includes a special case called the generalized noisy rare-weak model, which is defined as follows:
  • For any $j\in[p]$, $P_{\mathrm{clust}}^{(j)}$ is a generalized normal distribution with location parameter $x_{0}=0$, shape parameter $\gamma=\alpha$, and scale parameter $\sigma$.
  • The mean vector $\mu\in\mathbb{R}^{p}$ has only $s$ non-zero components, all of which are equal to $\lambda>0$, where $s$ and $\lambda$ are set appropriately to satisfy assumptions (A1)-(A4).
If we require the exponential Orlicz norm to be less than 1, we need to adjust the scale parameter $\sigma$ of $P_{\text{clust}}^{(j)}$. Donoho and Jin [11] studied this model where $\eta=0,\gamma=1$ and $\sigma=1$.
3 Main results
3.1 Generalization bound
We derive a generalization bound for the maximum margin classifier in a relaxed standard mixture model.
Theorem 4.
For any $\alpha\in(0,2]$ and $\kappa\in(0,1)$, there exists a constant $c>0$ such that, under assumptions (A1)-(A5), for all large enough $C$, with probability at least $1-\delta$, the maximum margin classifier $w$ satisfies
$\underset{(x,y)\sim P}{\mathbb{P}}\left[\mathrm{sign}(w\cdot x)\neq y\right]\leq\eta+\exp\left(-c\frac{\|\mu\|^{2\alpha}}{p^{\alpha/2}}\right).$
The proof of this theorem is provided in Section 4. This theorem reveals the relationship between the number of dimensions $p$ and $\|\mu\|$ in determining the success of learning. Specifically, when $\|\mu\|$ increases as $\|\mu\|=\Theta(p^{\tau})$ for any $\tau\in(1/4,1/2]$, the misclassification error of the maximum margin classifier asymptotically approaches the noise level $\eta$. The rate of increase in $\|\mu\|$ for benign overfitting is same as that proved by Chatterji and Long [7] when $\alpha=2$. Therefore, our result shows that in high-dimensional feature spaces, if the signal is sufficiently strong, learning can be achieved while minimizing the impact of noise even for heavier tails ($\alpha<2$).
Here are the implications of Theorem 4 in the noisy rare-weak model where the mean vector $\mu$ has only $s$ non-zero elements and all non-zero elements equal $\gamma$.
Corollary 5.
There exists a constant $c>0$ such that, under assumptions (A1)-(A5), in the generalized noisy rare-weak model, for any $\lambda\geq 0$ and all large enough $C$, with probability $1-\delta$, a maximum margin classifier $w$ satisfies
$\underset{(x,y)\sim P}{\mathbb{P}}\left[\mathrm{sign}(w\cdot x)\neq y\right]\leq\eta+\exp\left(-c\frac{(\lambda^{2}s)^{\alpha}}{p^{\alpha/2}}\right).$
We will consider $\lambda$ as fixed. Jin [15] demonstrated that for the noiseless rare-weak model, learning is impossible when $s=O(\sqrt{p})$ under the Gaussian assumption. Considering the fact that the Gaussian distribution is an $\alpha$ sub-gaussian for every $\alpha$ in $(0,2]$, their counterexample can show that our upper bound has optimality in a sense. Strictly speaking, to fit Jin’s model to our model, we need to adjust the scale parameter $\sigma$ of $P_{\text{clust}}^{(j)}$ to make the exponential Orlicz norm less than 1. However, this adjustment does not affect the accuracy of the maximum margin classifier.
3.2 Learning rate
We perform a detailed analysis of sufficient conditions for the learning rate when benign overfitting occurs. To concretely calculate the assumption of the learning rate, a bound of the largest singular value of $X$ is used.
Proposition 6 (A bound of the singular values of $X$).
For any $\delta\geq 0$, with probability at least $1-\delta$, there are constants $c_{5}$, $c_{6}$, $c_{7}$, $c_{8}$ depending only on $\alpha$ such that
$s_{1}(X)$
$\leq\sqrt{p}\Bigg{(}c_{5}+\frac{c_{6}\sqrt{n}}{p}\sum_{i=1}^{p}|\mu_{i}|+\frac{2n\|\mu\|^{2}}{p}$
$\quad+\frac{c_{7}+c_{8}\max_{i}|\mu_{i}|\sqrt{n}}{p}\left(n\log 9+\log\frac{4}{\delta}\right)^{\frac{2}{\alpha}}\Bigg{)}.$
The proof of Proposition 6 is in Appendix B.2. According to Proposition 6, a sufficient condition for assumption (A5) can be expressed as assumption (A6):
  1. (A6)
    The learning rate satisfies:
    $\beta\leq$
    $\min\Bigg{(}\frac{8}{p}\Bigg{(}c_{5}+\frac{c_{6}\sqrt{n}}{p}\sum_{i=1}^{p}|\mu_{i}|+\frac{2n\|\mu\|^{2}}{p}$
    $\quad+\frac{c_{7}+c_{8}\max_{i}|\mu_{i}|\sqrt{n}}{p}\left(n\log 9+\log\frac{4}{\delta}\right)^{\frac{2}{\alpha}}\Bigg{)}^{-2},$
    $\quad\frac{1}{c_{2}p}\left(1+\frac{2n}{p}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right)^{-1}\Bigg{)}.$
By using assumption (A6) instead of (A5), we obtain Corollary 7.
Corollary 7.
Under assumptions (A1)-(A4) and (A6) for all large enough $C$, with probability at least $1-2\delta$, the same generalization error bound as in Theorem 4 holds.
Moreover, by Corollary 7, we obtain Corollary 8 and 9. The proofs of Corollaries 8 and 9 are in Appendix B.3.
Corollary 8.
Under assumptions (A1)-(A4) for all large enough $C$, if $\beta$ satisfies
$\beta\leq c_{9}p^{-1}$
where $c_{9}$ is a constant depending on $\alpha,\kappa,\delta$, and $n$, with probability at least $1-2\delta$, the same generalization error bound as in Theorem 4 holds.
This corollary implies that when $p$ and $\|\mu\|$ grow large while $n$ and $\delta$ are fixed under assumptions (A3) and (A4), $\beta=O(p^{-1})$ is sufficient for the same result as Theorem 4. The order remains unchanged even when $\alpha$ is small.
Corollary 9.
Under assumptions (A1)-(A4) for all large enough $C$, if $\beta$ satisfies
$\beta\leq c_{10}p^{-1}\left(1+n^{\frac{2}{\alpha}-1}(\log n)^{-\frac{1}{\alpha}}\right)^{-2}$
where $c_{10}$ is a constant depending on $\alpha,\kappa$, and $\delta$, with probability at least $1-2\delta$, the same generalization error bound as in Theorem 4 holds.
Since $p\geq n^{2}$, it is straightforward to show that $\beta=O(p^{-\frac{2}{\alpha}})$ ensures the same generalization error bound as not only $p$ and $\|\mu\|$, but also $n$ grows under assumptions (A3) and (A4). As $\alpha$ decreases, the order also decreases, indicating that, for heavy-tailed distributions, the learning rate must be reduced accordingly.
4 Sketch of proof of Theorem 4
In the lemmas of this section, we assume (A1)-(A5). The proofs of the lemmas in this section are provided in Appendix B.1. For simplicity, we assume $U=I$. This assumption can be made without loss of generality for the following reasons:
  • Transformation of the maximum margin classifier:
    If $w$ is the maximum margin classifier for the original data points $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$, then $Uw$ becomes the maximum margin classifier for the transformed data points $(Ux_{1},y_{1}),\ldots,(Ux_{n},y_{n})$.
  • Probability equivalence:
    The probability of misclassification remains unchanged whether we consider $y(w\cdot x)<0$ or $y(Uw)\cdot(Ux)<0$.
For the same reason as in section 4 of [7], without loss of generality, we can assume $\mathbb{P}(x=\tilde{x})=1$ and $\mathbb{P}(y\neq\tilde{y})=\eta$.
We define the sets of indices of “noisy” and “clean” samples.
Definition 10.
Let $\mathcal{N}$ denote the set $\{k:y_{k}\neq\tilde{y}_{k}\}$ of indices of “noisy” samples, and $\mathcal{C}$ denote the set $\{k:y_{k}=\tilde{y}_{k}\}$ indices of “clean” samples.
Next, we define $z_{k}$, $\tilde{z}_{k}$, $\xi_{k}$, and $\tilde{\xi}_{k}$ to simplify the subsequent discussion.
Definition 11.
For index $k\in[n]$ of each example, let $z_{k}$ denote $x_{k}y_{k}$ and let $\tilde{z}_{k}$ denote $\tilde{x}_{k}\tilde{y}_{k}$. Let $\xi_{k}$ denote $z_{k}-\mathbb{E}[z_{k}]$ and let $\tilde{\xi}_{k}$ denote $\tilde{z}_{k}-\mathbb{E}[\tilde{z}_{k}]$.
Then, $\xi_{k}$ and $\tilde{\xi}_{k}$ are $\alpha$ sub-exponential, and the following lemma holds:
Lemma 12.
For any $k\in[n]$,
  1. 1.
    $\mathbb{E}[z_{k}]=\mathbb{E}[\tilde{z}_{k}]=\mu$ and
  2. 2.
    each component of $\xi_{k}$ and $\tilde{\xi}_{k}$ is $\alpha$ sub-exponential, with its exponential Orlicz norm at most $1$.
The next lemma provides an upper bound for the misclassification error. This bound is expressed in terms of two factors:
  • The expected value of the margin on unperturbed data points, denoted as
    $\underset{(\tilde{x},\tilde{y})\sim\tilde{P}}{\mathbb{E}}[\tilde{y}(w\cdot\tilde{x})],$
    which equals $w\cdot\mu$.
  • The Euclidean norm of the classifier vector $w$.
Lemma 13.
For any $w\in\mathbb{R}^{p}\setminus\{0\}$, there exists a positive constant $c$ such that
$\mathbb{P}_{(x,y)\sim P}\left[\mathrm{sign}(w\cdot x)\neq y\right]\leq\eta+2\exp\left(-c\frac{|w\cdot\mu|^{\alpha}}{\|w\|^{\alpha}}\right).$
The next lemma provides concentration arguments for $z_{k}$.
Lemma 14.
For any $\alpha\in(0,2]$ and $\kappa\in(0,1)$, there exists a constant $c_{1}\geq 1$ such that, for any $c^{\prime}$, for all large enough $C$, with probability at least $1-\delta$, the following holds:
  1. 1.
    For any $k\in[n]$,
    $\frac{p}{c_{1}}\leq\|z_{k}\|^{2}\leq c_{1}p.$
  2. 2.
    For any $i\neq j\in[n]$,
    $|z_{i}\cdot z_{j}|\leq c^{\prime}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right).$
  3. 3.
    For any $k\in\mathcal{C}$,
    $|\mu\cdot z_{k}-\|\mu\|^{2}|<\frac{\|\mu\|^{2}}{2}.$
  4. 4.
    For any $k\in\mathcal{N}$,
    $|\mu\cdot z_{k}-(-\|\mu\|^{2})|<\frac{\|\mu\|^{2}}{2}.$
  5. 5.
    The number of noisy samples satisfies $|\mathcal{N}|\leq(\eta+c^{\prime})n$.
  6. 6.
    The samples are linearly separable.
From here on, we will assume that samples satisfy all the conditions of Lemma 14.
The next lemma provides the bound on the ratio of losses when the loss function is the sigmoid loss.
Lemma 15.
There exists a positive constant $c_{3}$ such that, for all large enough $C$, and any learning rate $\beta$ which satisfies
$\beta\leq\frac{1}{2c_{1}}\left(p+2n\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right)^{-1},$
for all iterations $t\geq 0$,
$\max_{i,j\in[n]}\left\{\frac{1+\exp(\theta^{(t)}\cdot z_{j})}{1+\exp(\theta^{(t)}\cdot z_{i})}\right\}\leq c_{3},$
where $c_{1}$ is a constant which satisfies Lemma 14.
Soudry et al. [30] provide results regarding the convergence behavior of $\theta^{(t)}$ when the data is linearly separable.
Lemma 16 (Soudry et al., 2018 [30]).
For any linearly separable $\mathcal{S}$ and for $\beta\leq 8(s_{1}(X))^{-2}$, we have
$\frac{w}{\|w\|}=\lim_{t\rightarrow\infty}\frac{\theta^{(t)}}{\|\theta^{(t)}\|}.$
Using Lemmas 14, 15, and 16, we derive Lemma 17.
Lemma 17.
For any $\kappa\in(0,1)$, there exists a positive constant $c_{4}$ such that, for any large enough $C$, with probability at least $1-\delta$, the maximum margin weight vector $w$ satisfies,
$\mu\cdot w\geq\frac{\|w\|\|\mu\|^{2}}{c_{4}\sqrt{p}}.$
By Lemmas 13 and 17, we have Theorem 4.
5 Simulation
We conducted simulation studies to assess the performance of the maximum margin classifier across various conditions, specifically focusing on how dimensionality, tail heaviness, and learning rates interact. The simulation was designed with the following parameters: the training set consisted of $500$ samples, we used $1000$ test samples to assess generalization, and each experiment was repeated $5$ times, and the results were averaged to ensure robustness. The link to the detailed code for the experiments is provided in Appendix D.
Figure 3: A heatmap showing the mean test error for $\beta=0.001$ with the horizontal axis representing the dimension $p$ and the vertical axis representing the shape parameter $\gamma$.
Figure 4: A heatmap showing the mean test error for $\gamma=0.8$ with the horizontal axis representing the dimension $p$ and the vertical axis representing the learning rate $\beta$.
Figure 5: A heatmap showing the mean test error for $p=4000$ with the horizontal axis representing the shape parameter $\gamma$ and the vertical axis representing the learning rate $\beta$.
5.1 Data generation
The data was generated by the heavy-tailed setting, as described in Section 2.3. We set $P_{\mathrm{clust}}$ to be the generalized normal distribution with a scale parameter of $1$ and a shape parameter $\gamma$ ranging from $0.5$ to $1$. This distribution is used to control the tail behavior of the data, where smaller values of $\gamma$ correspond to heavier tails.
For the mean vector, $\mu$, the first $\lfloor p^{2/3}\rfloor$ elements were set to 1, and the remaining elements were set to 0, ensuring that $\|\mu\|=\Theta(p^{1/3})$. We chose an orthogonal matrix $U$ such that $U=I$, the identity matrix.
We also incorporated label noise by flipping the labels with a noise level of $\eta=0.05$, meaning that each true label was flipped with a probability of $\eta$.
5.2 Model training
We used the maximum margin classifier, as described in Section 2.4. The model was trained for $100000$ epochs to ensure convergence. We conducted three different numerical experiments to observe how these conditions influence test error.
5.3 Results and discussion
Experiment 1: interaction between dimension $p$ and shape parameter $\gamma$ (Figure 3)
In the first experiment, we investigated how the interaction between the dimensionality $p$ and the shape parameter $\gamma$ influences model performance. As $p$ increases, the test error initially decreases and then stabilizes around the noise level. For smaller values of $\gamma$ (heavier tails), the stabilization occurs more slowly, indicating that learning from heavy-tailed distributions is more challenging. In contrast, for larger values of $\gamma$ (lighter tails), the model converges faster.
These results suggest that high-dimensional parameter spaces permit benign overfitting, regardless of the tail heaviness. However, for heavier-tailed distributions (smaller $\gamma$), more dimensions are required to achieve similar performance compared to lighter-tailed distributions. This is consistent with the theoretical assumptions (A3) and (A4).
Experiment 2: impact of dimension $p$ and learning rate $\beta$ (Figure 4)
In the second experiment, we explored how the interaction between dimensionality $p$ and learning rate $\beta$ affects the test error. For large $p$, benign overfitting does not occur unless a small learning rate $\beta$ is chosen. If $\beta$ is too large, the learning process struggles to make progress.
As $p$ increases, the generalization error should decrease, as predicted by the benign overfitting bound. However, if the learning rate $\beta$ is not sufficiently small, the conditions outlined in Corollary 7 are not satisfied, and the learning process does not perform well. Our simulations suggest that in high-dimensional parameter spaces, the learning rate $\beta$ must be reduced to enable benign overfitting.
Experiment 3: impact of shape parameter $\gamma$ and learning rate $\beta$ (Figure 5)
In the third experiment, we fixed the dimensionality at $p=4000$ and examined the interaction between the shape parameter $\gamma$ and the learning rate $\beta$. For smaller $\gamma$ (heavier tails), the model is more sensitive to the choice of $\beta$. In particular, larger values of $\beta$ result in higher test errors for smaller $\gamma$. Conversely, for larger values of $\gamma$ (lighter tails), the model performs well even with larger learning rates. This suggests that careful tuning of the learning rate is crucial when dealing with heavy-tailed distributions to achieve benign overfitting.
These findings align with the theoretical results, indicating that when $\gamma$ is small, if $\beta$ is not sufficiently small, the condition on $\beta$ specified in Corollary 9 is violated.
6 Conclusion
Our research extends the analysis of benign overfitting in binary classification problems to heavy-tailed input distributions, specifically $\alpha$ sub-exponential distributions where $\alpha\in(0,2]$. The main findings of this study are:
  1. 1.
    We derived generalization bounds for maximum margin classifiers in this heavy-tailed setting, showing that benign overfitting can occur under certain conditions on dimensionality $p$ and the feature vector magnitude $\|\mu\|$.
  2. 2.
    Our results demonstrate that as the number of dimensions $p$ increases and the feature vector magnitude $\|\mu\|$ scales as $\Theta(p^{d})$ for $d\in(1/4,1/2]$, the misclassification error approaches the noise level $\eta$ even under the heavy-tailed setting.
  3. 3.
    In the context of the noisy rare-weak model, our upper bounds suggest that the maximum margin classifier can succeed arbitrarily close to the known impossibility threshold of $s=O(\sqrt{p})$.
  4. 4.
    We showed that the upper bound on the learning rate for benign overfitting, and demonstrated that when $n$ is fixed, the bound is of order $p^{-1}$, while in the case where $n,p,\|\mu\|$ are large, we observed that the upper bound decreases as $\alpha$ increases.
  5. 5.
    By conducting simulations, we confirmed that the relationship between the number of parameters, the tail heaviness, and the learning rate when benign overfitting occurs follows the same trend as that derived theoretically.
These findings significantly contribute to our understanding of benign overfitting by showing that the phenomenon is not limited to sub-gaussian distributions but extends to heavier-tailed inputs as well. This research bridges a gap between theory and practice, as real-world data often exhibit heavier tails than the Gaussian distribution.
Our work opens up several avenues for future research:
  1. 1.
    Investigation of benign overfitting in even heavier-tailed distributions, such as those with polynomial tails.
  2. 2.
    Extension of the analysis to multi-class classification problems with heavy-tailed inputs.
  3. 3.
    Exploration of the implications of these findings for deep learning models, which often deal with high-dimensional, heavy-tailed data.
  4. 4.
    Development of new learning algorithms that explicitly leverage the properties of heavy-tailed distributions to achieve better generalization in high-dimensional settings.
In conclusion, this study provides a significant step towards understanding the phenomenon of benign overfitting in more realistic data settings. By extending the theory to heavy-tailed distributions, we have broadened the applicability of benign overfitting results to a wider range of practical scenarios, potentially impacting the design and analysis of machine learning algorithms for complex, real-world data.
Acknowledgements
We would like to thank our colleagues for their valuable feedback and suggestions that helped improve this work. This research was supported in part by RIKEN AIP and JSPS KAKENHI (JP22K03439).
Supplementary Materials
Appendix A Concentration inequality
In this section, we introduce the concentration inequalities for $\alpha$ sub-exponential random variables. In our proof, we apply the following two concentration inequalities.
Proposition 18 (A special case of Theorem 1.5 in [13]).
Let $\alpha\in(0,2]$ and $K$ be a positive constant and $a\in\mathbb{R}^{n}$ be a constant vector. Let $X_{1},\ldots,X_{n}$ be independent mean-zero random variables satisfying $\|X_{i}\|_{\psi_{\alpha}}\leq K$. Then, there exists a positive constant $c_{\alpha}$ such that for any $t>0$ it holds
$\mathbb{P}\left[\left|\sum_{i=1}^{n}a_{i}X_{i}\right|\geq t\right]\leq 2\exp\left(-\frac{1}{c_{\alpha}}\frac{t^{\alpha}}{K^{\alpha}\|a\|^{\alpha}}\right).$
This is a special case of Theorem 1.5 of Götze et al. [13].
Theorem 19 (Extended Hanson-Wright inequality [29]).
Let $\alpha\in(0,2]$ and $K$ be a positive constant. Let $X_{1},\ldots,X_{n}$ be independent mean-zero random variables such that $\|X_{i}\|_{\psi_{\alpha}}\leq K$, the corresponding random vector $X$ be $(X_{1},\ldots,X_{n})^{T}$ and $A\in\mathbb{R}^{n\times n}$ be a symmetric matrix. Then, there exists a positive constant $c_{\alpha}$ such that it holds
$\mathbb{P}[|X^{T}AX-\mathbb{E}[X^{T}AX]|\geq t]\leq 2\exp\left(-\frac{1}{c_{\alpha}}\min\left(\frac{t^{2}}{K^{4}\|A\|_{\mathrm{HS}}^{2}},\left(\frac{t}{K^{2}\|A\|_{\mathrm{op}}}\right)^{\frac{\alpha}{2}}\right)\right)$
for any $t\geq 0$.
Appendix B Missing proofs
B.1 Proofs of lemmas used in the proof of Theorem 4
B.1.1 Proof of Lemma 12
Proof of Lemma 12.
By definition of $\tilde{z}_{k}$ and $z_{k}$,
$\mathbb{E}[\tilde{z}_{k}]=\mathbb{E}[\tilde{x}_{k}\tilde{y}_{k}]=\mathbb{E}[(q+\tilde{y}_{k}\mu)\tilde{y}_{k}]=\mu,$
$\mathbb{E}[z_{k}]=\mathbb{E}[\tilde{x}_{k}y_{k}]=\mathbb{E}[(q+y_{k}\mu)y_{k}]=\mu.$
Together with the definition of $\xi_{k}$,
$\|\xi_{kl}\|_{\psi_{\alpha}}=\left\|q_{l}y_{k}\right\|_{\psi_{\alpha}}=\left\|q_{l}\right\|_{\psi_{\alpha}}\leq 1,$
$\|\tilde{\xi}_{kl}\|_{\psi_{\alpha}}=\left\|q_{l}\tilde{y}_{k}\right\|_{\psi_{\alpha}}=\left\|q_{l}\right\|_{\psi_{\alpha}}\leq 1.$
B.1.2 Proof of Lemma 13
Proof of Lemma 13.
Following the proof of Chatterji and Long [7], we have
$\mathbb{P}_{(x,y)\sim P}\left[\mathrm{sign}(w\cdot x)\neq y\right]$ $=\mathbb{P}_{(x,y)\sim P}\left[y(w\cdot x)<0\right]$
$\leq\eta+\mathbb{P}_{(\tilde{x},\tilde{y})\sim\tilde{P}}\left[\tilde{y}(w\cdot\tilde{x})<0\right]$
$=\eta+\mathbb{P}_{(\tilde{x},\tilde{y})\sim\tilde{P}}\left[\left(\frac{w}{\|w\|}\cdot\tilde{y}\tilde{x}\right)<0\right]$
$=\eta+\mathbb{P}_{(\tilde{x},\tilde{y})\sim\tilde{P}}\left[\left(\frac{w}{\|w\|}\cdot\tilde{\xi}\right)<-\frac{w}{\|w\|}\cdot\mu\right].$
Applying Proposition 18, there exists a positive constant $c$ such that
$\mathbb{P}_{(\tilde{x},\tilde{y})\sim\tilde{P}}\left[\left(\frac{w}{\|w\|}\cdot\tilde{\xi}\right)<-\frac{w}{\|w\|}\cdot\mu\right]$ $\leq 2\exp\left(-c\frac{|w\cdot\mu|^{\alpha}}{\|w\|^{\alpha}}\right),$
which completes our proof. ∎
B.1.3 Proof of Lemma 14
In this section, we prove Lemma 14 by using concentration inequalities from Section A. We assume assumptions (A1)-(A4) hold, and decompose Lemma 14 into six different parts. We prove that each separate lemma holds with probability at least $61-\delta/$ 6.
Lemma 20.
For any $\alpha\in(0,2]$ and $\kappa\in(0,1)$, there exists a constant $c\geq 1$ such that, for all large enough $C$, with probability at least $61-\delta/$ 6, for any $k\in[n]$,
$\frac{p}{c}\leq\|z_{k}\|^{2}\leq cp.$
Proof.
By Theorem 19 with $A=I$, there exists a positive constant $c$ such that
$\mathbb{P}\left[\left|\|\xi_{k}\|^{2}-\mathbb{E}[\|\xi_{k}\|^{2}]\right|\geq t\right]\leq 2\exp\left(-\frac{1}{c}\min\left(\frac{t^{2}}{p},t^{\frac{\alpha}{2}}\right)\right).$
By setting $t=\frac{\kappa p}{2}$
$2\exp\left(-\frac{1}{c}\min\left(\frac{t^{2}}{p},t^{\frac{\alpha}{2}}\right)\right)=2\exp\left(-\frac{1}{c}\min\left(\left(\frac{\kappa}{2}\right)^{2}p,\left(\frac{\kappa}{2}\right)^{\frac{\alpha}{2}}p^{\frac{\alpha}{2}}\right)\right).$
By assumption (A3), we have $p\geq C\left(\log\frac{n}{\delta}\right)^{\frac{2}{\alpha}}$. There exists a large enough constant $C$ such that
$2\exp\left(-\frac{1}{c}\min\left(\left(\frac{\kappa}{2}\right)^{2}p,\left(\frac{\kappa}{2}\right)^{\frac{\alpha}{2}}p^{\frac{\alpha}{2}}\right)\right)\leq\frac{\delta}{6n}.$
Thus,
$\mathbb{P}\left[\left|\|\xi_{k}\|^{2}-\mathbb{E}[\|\xi_{k}\|^{2}]\right|\geq\frac{\kappa p}{2}\right]\leq\frac{\delta}{6n}.$ (1)
Recalling the assumption $\mathbb{E}[\|q\|^{2}]\geq\kappa p$, we have
$\mathbb{E}[\|\xi_{k}\|^{2}]=\mathbb{E}\left[\|z_{k}-\mathbb{E}[z]\|^{2}\right]=\mathbb{E}[\|q\|^{2}]\geq\kappa p.$
Let $\{\xi_{kj}\}_{j=1}^{p}$ be elements of $\xi_{k}$. By $\|\xi_{kj}\|_{\psi_{\alpha}}\leq 1$ for each $j$,
[…]
$=\frac{4}{\alpha}\Gamma\left(\frac{2}{\alpha}\right).$
Thus, $\mathbb{E}[\|\xi_{k}\|^{2}]\leq\frac{4p}{\alpha}\Gamma\left(\frac{2}{\alpha}\right)$. Because of this and (1), with probability at least $1-\frac{\delta}{6n}$,
$\frac{\kappa p}{2}\leq\|\xi_{k}\|^{2}\leq\left(\frac{4}{\alpha}\Gamma\left(\frac{2}{\alpha}\right)+\frac{\kappa}{2}\right)p.$
Suppose $k\in\mathcal{C}$, and let $\xi_{k}=z_{k}-\mu$. Then, the following inequalities hold.
  1. 1.
    $\|z_{k}-\mu\|^{2}\leq 2\|z_{k}\|^{2}+2\|\mu\|^{2}$.
  2. 2.
    $\|\mu\|^{2}<\frac{p}{C}$ by assumption (A3).
  3. 3.
    $\|\xi_{k}\|^{2}=\|z_{k}-\mu\|^{2}\geq\frac{\kappa p}{2}$ with probability at least $1-\frac{\delta}{6}$.
Combining these inequalities, we obtain
$\|z_{k}\|^{2}$ $\geq\frac{1}{2}\|z_{k}-\mu\|^{2}-\|\mu\|^{2}$
$\geq\frac{1}{2}\left(\frac{\kappa p}{2}\right)-\left(\frac{p}{C}\right)$
$=\frac{\kappa p}{4}-\frac{p}{C}.$
For sufficiently large $C$, we can ensure that $\frac{p}{C}<\frac{\kappa p}{8}$. Thus,
$\|z_{k}\|^{2}>\frac{\kappa p}{4}-\frac{\kappa p}{8}=\frac{\kappa p}{8}.$ (2)
Therefore, with probability at least $1-\frac{\delta}{6n}$, we have $\|z_{k}\|^{2}>\frac{\kappa p}{8}$ for sufficiently large $C$. Again by $\|z_{k}\|^{2}\leq 2\|z_{k}-\mu\|^{2}+2\|\mu\|^{2}$,
$\|z_{k}\|^{2}$ $\leq 2\|z_{k}-\mu\|^{2}+2\|\mu\|^{2}$
$\leq 2\left(\frac{4}{\alpha}\Gamma\left(\frac{2}{\alpha}\right)+\frac{\kappa}{2}\right)p+2\|\mu\|^{2}$
$<2\left(\frac{4}{\alpha}\Gamma\left(\frac{2}{\alpha}\right)+\frac{\kappa}{2}\right)p+\frac{2p}{C}$
$<\left(\frac{8}{\alpha}\Gamma\left(\frac{2}{\alpha}\right)+\kappa+2\right)p.$
Therefore, setting $c=\max\left(\frac{8}{\kappa},\frac{8}{\alpha}\Gamma\left(\frac{2}{\alpha}\right)+\kappa+2\right)$, we have
$\frac{p}{c}\leq\|z_{k}\|^{2}\leq cp$
for any $k\in\mathcal{C}$. A similar argument holds for $k\in\mathcal{N}$. ∎
Lemma 21.
There exists $c\geq 1$ such that, for any large enough $C$, with probability at least $1-\frac{\delta}{6}$, for any $i\neq j\in[n]$,
$|z_{i}\cdot z_{j}|\leq c\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right).$
Proof.
Applying Theorem 19 as in the proof of Lemma 20 and using the union bound method, we have
$\mathbb{P}[\exists i\in[n],\|\xi_{i}\|\geq\sqrt{p}]\leq\frac{\delta}{24}.$
For any pair $i,j\in[n]$, we have
$\mathbb{P}[|\xi_{i}\cdot\xi_{j}|\geq t]\leq\mathbb{P}[|\xi_{i}\cdot\xi_{j}|\geq t\mid\|\xi_{j}\|\leq\sqrt{p}]+\mathbb{P}[\|\xi_{j}\|>\sqrt{p}].$ (3)
Regarding $\xi_{j}$ as fixed, by Proposition 18 there exists a positive constant $c_{2}$ such that
$\mathbb{P}[|\xi_{i}\cdot\xi_{j}|\geq t]=\mathbb{P}\left[\frac{\xi_{j}}{\|\xi_{j}\|}\cdot\xi_{i}\geq\frac{t}{\|\xi_{j}\|}\right]\leq 2\exp\left(-c_{2}\frac{t^{\alpha}}{\|\xi_{j}\|^{\alpha}}\right).$
Therefore,
$\mathbb{P}[|\xi_{i}\cdot\xi_{j}|\geq t\mid\|\xi_{j}\|\leq\sqrt{p}]$ $\leq 2\exp\left(-c_{2}\frac{t^{\alpha}}{p^{\frac{\alpha}{2}}}\right),$
$\mathbb{P}[|\xi_{i}\cdot\xi_{j}|\geq t]$ $\leq 2\exp\left(-c_{2}\frac{t^{\alpha}}{p^{\frac{\alpha}{2}}}\right)+\mathbb{P}[\|\xi_{j}\|>\sqrt{p}].$
By the union bound method,
$\mathbb{P}\left[\exists i\neq j\in[n],|\xi_{i}\cdot\xi_{j}|\geq t\right]\leq 2n^{2}\exp\left(-c_{2}\frac{t^{\alpha}}{p^{\frac{\alpha}{2}}}\right)+\mathbb{P}[\exists j\in[n],\|\xi_{j}\|>\sqrt{p}].$
Setting $t=c_{3}\left(p^{\frac{\alpha}{2}}\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}$, for a large enough $c_{3}$, we have
$\mathbb{P}\left[\exists i\neq j\in[n],|\xi_{i}\cdot\xi_{j}|\geq c_{3}\left(p^{\frac{\alpha}{2}}\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right]\leq\frac{\delta}{24}+\mathbb{P}[\exists j\in[n],\|\xi_{j}\|>\sqrt{p}].$
Together with the inequality (3), we have,
$\mathbb{P}\left[\exists i\neq j\in[n],|\xi_{i}\cdot\xi_{j}|\geq c_{3}\left(p^{\frac{\alpha}{2}}\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right]\leq\frac{\delta}{12}.$ (4)
By Proposition 18, there exists a constant $c_{4}$ such that
$\mathbb{P}[|\mu\cdot z_{k}|>\|\mu\|^{2}]$ $=\mathbb{P}\left[\left|\frac{\mu}{\|\mu\|}\cdot z_{k}\right|>\|\mu\|\right]\leq 2\exp\left(-c_{4}\|\mu\|^{\alpha}\right).$
By assumption (A4), for large enough $C$ we have
$\mathbb{P}[|\mu\cdot z_{k}|>\|\mu\|^{2}]\leq\frac{\delta}{12n}.$
By taking a union bound, we have
$\mathbb{P}[\exists k,|\mu\cdot z_{k}|>\|\mu\|^{2}]\leq\frac{\delta}{12}.$ (5)
Due to inequalities (4) and (5), with probability at least $1-\frac{\delta}{6}$,
$|z_{i}\cdot z_{j}|$ $=\Big{|}(z_{i}-\mathbb{E}[z_{i}])\cdot(z_{j}-\mathbb{E}[z_{j}])-\mathbb{E}[z_{i}]\cdot\mathbb{E}[z_{j}]+\mathbb{E}[z_{i}]\cdot z_{j}+\mathbb{E}[z_{j}]\cdot z_{i}\Big{|}$
$=\left|\xi_{i}\cdot\xi_{j}-\|\mu\|^{2}+\mu\cdot z_{j}+\mu\cdot z_{i}\right|$
$\leq|\xi_{i}\cdot\xi_{j}|+\|\mu\|^{2}+|\mu\cdot z_{j}|+|\mu\cdot z_{i}|$
$\leq 3\|\mu\|^{2}+c\left(p^{\frac{\alpha}{2}}\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}.$
Lemma 22.
For any large enough $C$, with probability at least $1-\frac{\delta}{6}$, for $k\in\mathcal{C}$,
$|\mu\cdot z_{k}-\|\mu\|^{2}|\leq\frac{\|\mu\|^{2}}{2}.$
Lemma 23.
For any large enough $C$, with probability at least $1-\frac{\delta}{6}$, for $k\in\mathcal{N}$,
$|\mu\cdot z_{k}-(-\|\mu\|^{2})|\leq\frac{\|\mu\|^{2}}{2}.$
Lemma 22 and Lemma 23 can be proven by the same logic. We will prove Lemma 22.
Proof.
If $k\in\mathcal{C}$,
$\mu\cdot z_{k}-\|\mu\|^{2}=\mu\cdot\xi_{k}.$
Since the exponential Orlicz norm of $\xi_{k}$ is at most $1$, by Proposition 18, there exists a positive constant $c$ such that
$\mathbb{P}\left[\left|\mu\cdot z_{k}-\|\mu\|^{2}\right|\geq\frac{\|\mu\|^{2}}{2}\right]$ $=\mathbb{P}\left[\left|\frac{\mu}{\|\mu\|}\cdot\xi_{k}\right|\geq\frac{\|\mu\|}{2}\right]$
$\leq 2\exp\left(-c\frac{\|\mu\|^{\alpha}}{2^{\alpha}}\right).$
By assumption (A4), for large enough constant $C$ we have
$\mathbb{P}\left[\left|\mu\cdot z_{k}-\|\mu\|^{2}\right|\geq\frac{\|\mu\|^{2}}{2}\right]\leq\frac{\delta}{6n}.$
Taking a union bound, we have
$\mathbb{P}\left[\exists k\in\mathcal{C},\left|\mu\cdot z_{k}-\|\mu\|^{2}\right|\geq\frac{\|\mu\|^{2}}{2}\right]\leq\frac{\delta}{6},$
which completes our proof. ∎
Lemma 24.
For any $c^{\prime}>0$, for any large enough $C$, with probability at least $1-\frac{\delta}{6}$, the number of noisy samples satisfies $|\mathcal{N}|\leq(\eta+c^{\prime})n$.
Proof.
$\mathbb{E}\left[|\mathcal{N}|\right]=\sum_{k=1}^{n}\mathbb{E}\left[1_{\{y_{k}\neq\tilde{y}_{k}\}}\right]=\sum_{k=1}^{n}\mathbb{P}[y_{k}\neq\tilde{y}_{k}]=n\eta.$
By Hoeffding’s inequality,
$\mathbb{P}\left[|\mathcal{N}|\geq(\eta+c^{\prime})n\right]$ $=\mathbb{P}\left[\frac{1}{n}\sum_{k=1}^{n}\left(1_{\{y_{k}\neq\tilde{y}_{k}\}}-\mathbb{E}[1_{\{y_{k}\neq\tilde{y}_{k}\}}]\right)\geq c^{\prime}\right]$
$\leq 2\exp\left(-2{c^{\prime}}^{2}n\right)$
$\leq\frac{\delta}{6}.$
The last inequality holds due to assumption (A2). ∎
Lemma 25.
If the following conditions hold, for any large enough $C$, $\{(x_{k},y_{k})\}_{k=1}^{n}$ are linearly separable.
  1. 1.
    There exists a positive constant $c$ such that for any $k\in[n]$
    $\frac{p}{c}\leq\|z_{k}\|^{2}\leq cp.$
  2. 2.
    There exists a positive constant $c$ such that for any $i\neq j\in[n]$
    $|z_{i}\cdot z_{j}|\leq c\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right).$
Proof.
Let $v$ be $\sum_{i\in[n]}z_{i}$. For each $k\in[n]$ and any $\delta>0$,
$y_{k}v\cdot x_{k}$ $=\sum_{i\in[n]}z_{i}\cdot z_{k}$
$=\|z_{k}\|^{2}+\sum_{i\neq k}z_{i}\cdot z_{k}$
$\geq\|z_{k}\|^{2}-\sum_{i\neq k}|z_{i}\cdot z_{k}|$
$\geq\frac{p}{c}-cn\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)$
$\geq\frac{p}{c}-2cn\max\left(\|\mu\|^{2},\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)$
$=\frac{1}{c}\left(p-2c^{2}n\max\left(\|\mu\|^{2},\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right).$
By assumptions (A3) and (A4), for large enough $C$ we have
$y_{k}v\cdot x_{k}>0,$
which completes our proof. ∎
B.1.4 Proof of Lemma 15
In this section, we will assume that samples satisfy all the conditions of Lemma 14. First, we will prove that the ratio of the losses between any pair of points is bounded. In this proof, we use Lemma 26, and Lemma 15 is derived from Lemma 26 and Lemma 27.
Lemma 26.
For any $s_{1},s_{2}\in\mathbb{R}$,
$\frac{1+\exp(s_{2})}{1+\exp(s_{1})}\leq\max\left(2,2\frac{\exp(-s_{1})}{\exp(-s_{2})}\right).$
Lemma 27.
There exists a positive constant $c_{2}$ such that, for all large enough $C$, and any learning rate $\beta$ which satisfies
$\beta\leq\frac{1}{2}\left(c_{1}p+2nc_{1}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right)^{-1},$
for all iterations $t\geq 0$,
$\max_{i,j\in[n]}\left\{\frac{\exp(-\theta^{(t)}\cdot z_{i})}{\exp(-\theta^{(t)}\cdot z_{j})}\right\}\leq c_{2},$
where $c_{1}$ is a constant which satisfies Lemma 14.
Proof.
For simplicity, let $A_{t}$ be the ratio between exponential losses of the first and second samples for $t$ iterations:
$A_{t}=\frac{\exp(-\theta^{(t)}\cdot z_{1})}{\exp(-\theta^{(t)}\cdot z_{2})}.$
We will show that $A_{t}\leq 4c_{1}^{2}$ by using induction. When $t=0$, $A_{0}=1\leq 4c_{1}^{2}$. Thus, the base step holds. Assume that the inductive hypothesis holds for some iteration $t$. We shall now show that it must hold at iteration $t+1$.
$A_{t+1}$ $=\frac{\exp(-\theta^{(t+1)}\cdot z_{1})}{\exp(-\theta^{(t+1)}\cdot z_{2})}$
$=\frac{\exp\left(-(\theta^{(t)}-\beta\nabla R(\theta^{(t)}))\cdot z_{1}\right)}{\exp\left(-(\theta^{(t)}-\beta\nabla R(\theta^{(t)}))\cdot z_{2}\right)}$
$=A_{t}\frac{\exp\left(\beta\nabla R(\theta^{(t)})\cdot z_{1}\right)}{\exp\left(\beta\nabla R(\theta^{(t)})\cdot z_{2}\right)}$
$=A_{t}\frac{\exp\left(-\beta\sum_{k\in[n]}\frac{z_{k}\cdot z_{1}}{1+\exp(\theta^{(t)}\cdot z_{k})}\right)}{\exp\left(-\beta\sum_{k\in[n]}\frac{z_{k}\cdot z_{2}}{1+\exp(\theta^{(t)}\cdot z_{k})}\right)}$
$=A_{t}\frac{\exp\left(-\beta\frac{\|z_{1}\|^{2}}{1+\exp(\theta^{(t)}\cdot z_{1})}\right)}{\exp\left(-\beta\frac{\|z_{2}\|^{2}}{1+\exp(\theta^{(t)}\cdot z_{2})}\right)}\frac{\exp\left(-\beta\sum_{k\neq 1}\frac{z_{k}\cdot z_{1}}{1+\exp(\theta^{(t)}\cdot z_{k})}\right)}{\exp\left(-\beta\sum_{k\neq 2}\frac{z_{k}\cdot z_{2}}{1+\exp(\theta^{(t)}\cdot z_{k})}\right)}$
$=A_{t}\exp\left(-\beta\left(\frac{\|z_{1}\|^{2}}{1+\exp(\theta^{(t)}\cdot z_{1})}-\frac{\|z_{2}\|^{2}}{1+\exp(\theta^{(t)}\cdot z_{2})}\right)\right)$
×$\quad\times\exp\left(-\beta\left(\sum_{k\neq 1}\frac{z_{k}\cdot z_{1}}{1+\exp(\theta^{(t)}\cdot z_{k})}-\sum_{k\neq 2}\frac{z_{k}\cdot z_{2}}{1+\exp(\theta^{(t)}\cdot z_{k})}\right)\right).$
By Lemma 14, for any $k,i\neq j\in[n]$, there exists a constant $c_{1}$ such that
$\frac{p}{c_{1}}\leq\|z_{k}\|^{2}\leq c_{1}p,$
$|z_{i}\cdot z_{j}|\leq c_{1}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right).$
Thus,
$A_{t+1}$ $\leq A_{t}\exp\left(-\beta\left(\frac{p/c_{1}}{1+\exp(\theta^{(t)}\cdot z_{1})}-\frac{c_{1}p}{1+\exp(\theta^{(t)}\cdot z_{2})}\right)\right)$
×$\quad\times\exp\left(2\beta\sum_{k\in[n]}\frac{c_{1}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)}{1+\exp(\theta^{(t)}\cdot z_{k})}\right)$
$=A_{t}\exp\left(-\frac{\beta p}{c_{1}(1+\exp(\theta^{(t)}\cdot z_{2}))}\left(\frac{1+\exp(\theta^{(t)}\cdot z_{2})}{1+\exp(\theta^{(t)}\cdot z_{1})}-c_{1}^{2}\right)\right)$
×$\quad\times\exp\left(2\beta\sum_{k\in[n]}\frac{c_{1}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)}{1+\exp(\theta^{(t)}\cdot z_{k})}\right).$
Now we consider two disjoint cases.
Case 1 ($A_{t}<2c_{1}^{2}$):
$A_{t+1}$ $\leq A_{t}\exp\left(\frac{\beta c_{1}p}{1+\exp(\theta^{(t)}\cdot z_{2})}\right)\times\exp\left(2\beta nc_{1}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right)$
$\leq A_{t}\exp\left(\beta\left(c_{1}p+2nc_{1}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right)\right).$
Taking $\beta$ small enough that
$\beta\leq\frac{1}{2}\left(c_{1}p+2nc_{1}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right)^{-1},$
we have
$A_{t+1}\leq A_{t}\exp\left(\frac{1}{2}\right)\leq 2c_{1}^{2}\exp\left(\frac{1}{2}\right)<4c_{1}^{2}.$
Case 2 ($A_{t}\geq 2c_{1}^{2}$):
$A_{t+1}$ $=A_{t}\exp\left(-\frac{\beta p}{c_{1}(1+\exp(\theta^{(t)}\cdot z_{2}))}\left(\frac{1+\exp(\theta^{(t)}\cdot z_{2})}{1+\exp(\theta^{(t)}\cdot z_{1})}-c_{1}^{2}\right)\right)$
×$\quad\times\exp\left(2\beta c_{1}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\frac{1}{1+\exp(\theta^{(t)}\cdot z_{2})}\sum_{k\in[n]}\frac{1+\exp(\theta^{(t)}\cdot z_{2})}{1+\exp(\theta^{(t)}\cdot z_{k})}\right).$
By Lemma 26 and the induction hypothesis,
$A_{t+1}$ $\leq A_{t}\exp\left(-\frac{\beta c_{1}p}{1+\exp(\theta^{(t)}\cdot z_{2})}\right)$
×$\quad\times\exp\left(2\beta c_{1}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\frac{1}{1+\exp(\theta^{(t)}\cdot z_{2})}\sum_{k\in[n]}\max\left(2,2A_{t}\right)\right)$
$\leq A_{t}\exp\left(-\frac{\beta c_{1}p}{1+\exp(\theta^{(t)}\cdot z_{2})}\right)$
×$\quad\times\exp\left(2\beta c_{1}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\frac{n\max\left(2,8c_{1}^{2}\right)}{1+\exp(\theta^{(t)}\cdot z_{2})}\right)$
$\leq A_{t}\exp\left(-\frac{\beta c_{1}}{1+\exp(\theta^{(t)}\cdot z_{2})}\left(p-8c_{1}^{2}n\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right)\right).$
By assumptions (A3) and (A4), for large enough $C$,
$p-8c_{1}^{2}n\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)>0.$
Thus,
$A_{t+1}<A_{t}\leq 4c_{1}^{2}.$
This completes the proof of the inductive step. ∎
B.1.5 Proof of Lemma 17
Proof of Lemma 17.
$\mu\cdot\theta^{(t+1)}$ $=\mu\cdot\theta^{(t)}+\beta\sum_{k\in[n]}\frac{\mu\cdot z_{k}}{1+\exp(\theta^{(t)}\cdot z_{k})}$
$=\mu\cdot\theta^{(t)}+\beta\sum_{k\in\mathcal{C}}\frac{\mu\cdot z_{k}}{1+\exp(\theta^{(t)}\cdot z_{k})}+\beta\sum_{k\in\mathcal{N}}\frac{\mu\cdot z_{k}}{1+\exp(\theta^{(t)}\cdot z_{k})}.$
By Lemma 14,
$\mu\cdot\theta^{(t+1)}$ $\geq\mu\cdot\theta^{(t)}+\frac{\beta\|\mu\|^{2}}{2}\sum_{k\in\mathcal{C}}\frac{1}{1+\exp(\theta^{(t)}\cdot z_{k})}-\frac{3\beta\|\mu\|^{2}}{2}\sum_{k\in\mathcal{N}}\frac{1}{1+\exp(\theta^{(t)}\cdot z_{k})}$
$\geq\mu\cdot\theta^{(t)}+\frac{\beta\|\mu\|^{2}}{2}\sum_{k\in[n]}\frac{1}{1+\exp(\theta^{(t)}\cdot z_{k})}-2\beta\|\mu\|^{2}\sum_{k\in\mathcal{N}}\frac{1}{1+\exp(\theta^{(t)}\cdot z_{k})}.$
By $|\mathcal{N}|\leq(\eta+c^{\prime})n$ and Lemma 15,
$\sum_{k\in\mathcal{N}}\frac{1}{1+\exp(\theta^{(t)}\cdot z_{k})}$ $\leq c_{3}(\eta+c^{\prime})n\min_{k\in[n]}\frac{1}{1+\exp(\theta^{(t)}\cdot z_{k})}$
$\leq c_{3}(\eta+c^{\prime})\sum_{k\in[n]}\frac{1}{1+\exp(\theta^{(t)}\cdot z_{k})},$
where $c_{3}$ is the constant from Lemma 15. Recalling that $\eta\leq\frac{1}{C}$ and $c^{\prime}$ is an arbitrary positive constant, for large enough $C$ and small enough $c^{\prime}$,
$\sum_{k\in\mathcal{N}}\frac{1}{1+\exp(\theta^{(t)}\cdot z_{k})}\leq\frac{1}{8}\sum_{k\in[n]}\frac{1}{1+\exp(\theta^{(t)}\cdot z_{k})}.$
Thus, we have
$\mu\cdot\theta^{(t+1)}$ $\geq\mu\cdot\theta^{(t)}+\frac{\beta\|\mu\|^{2}}{4}\sum_{k\in[n]}\frac{1}{1+\exp(\theta^{(t)}\cdot z_{k})}.$
By using this inequality repeatedly and $\theta^{(0)}=0$,
$\mu\cdot\theta^{(t+1)}$ $\geq\frac{\beta\|\mu\|^{2}}{4}\sum_{m=0}^{t}\sum_{k\in[n]}\frac{1}{1+\exp(\theta^{(m)}\cdot z_{k})},$
$\|w\|\frac{\mu\cdot\theta^{(t+1)}}{\|\theta^{(t+1)}\|}$ $\geq\|w\|\frac{\beta\|\mu\|^{2}\sum_{m=0}^{t}\sum_{k\in[n]}\frac{1}{1+\exp(\theta^{(m)}\cdot z_{k})}}{4\|\theta^{(t+1)}\|}.$
By taking the large-$t$ limit and using Lemma 16,
$\mu\cdot w\geq\beta\|w\|\|\mu\|^{2}\lim_{t\rightarrow\infty}\frac{\sum_{m=0}^{t}\sum_{k\in[n]}\frac{1}{1+\exp(\theta^{(m)}\cdot z_{k})}}{4\|\theta^{(t+1)}\|}$ (6)
By definition of gradient descent iterations,
$\|\theta^{(t+1)}\|$ $=\left\|\sum_{m=0}^{t}\beta\nabla R(\theta^{(m)})\right\|$
$\leq\beta\sum_{m=0}^{t}\|\nabla R(\theta^{(m)})\|$
$\leq\beta\sum_{m=0}^{t}\left\|\sum_{k\in[n]}\frac{-z_{k}}{1+\exp(\theta^{(m)}\cdot z_{k})}\right\|$
$\leq\beta c_{1}\sqrt{p}\sum_{m=0}^{t}\sum_{k\in[n]}\frac{1}{1+\exp(\theta^{(m)}\cdot z_{k})}.$
With inequality (6), we have
$\mu\cdot w\geq\frac{\|w\|\|\mu\|^{2}}{4c_{1}\sqrt{p}},$
which completes our proof. ∎
B.2 Proof of Proposition 6
Let $\tilde{X}$ denote $[\tilde{y}_{1}\tilde{x}_{1},\cdots,\tilde{y}_{n}\tilde{x}_{n}]\in\mathbb{R}^{p\times n}$ where $\tilde{x}_{k}=q_{k}+\mu\tilde{y}_{k}$. $\{q_{k}\}$ and $\{y_{k}\}$ are independent of each other. $q_{k}\sim P_{\text{clust}}$ and $\tilde{y}_{k}\sim\text{Uniform}\{-1,1\}$ for each $k\in[n]$. Let $\tilde{y}=[\tilde{y}_{1},\dots,\tilde{y}_{n}]^{T}$. Before proving Proposition 6, we first present Proposition 28 for a simpler case.
Proposition 28 (A bound of the singular values of $\tilde{X}$).
We assume $\sigma^{2}=\mathbb{E}_{q_{i}\sim P_{\text{clust}}^{(i)}}[q_{i}^{2}]$ for any $i\in[n]$. For any $\delta\geq 0$, with probability at least $1-\delta$, there are constants $c_{3}$, $c_{4}$ depending on $\alpha$ such that
$s_{1}(\tilde{X})\leq\sigma\sqrt{p}\left(1+\frac{2n\|\mu\|^{2}}{\sigma^{2}p}+\frac{c_{3}+c_{4}\max_{i}|\mu_{i}|\sqrt{n}}{\sigma^{2}p}\left(n\log 9+\log\frac{4}{\delta}\right)^{\frac{2}{\alpha}}\right).$
This bound also holds for $X=[y_{1}\tilde{x}_{1},\cdots,y_{n}\tilde{x}_{n}]$, which consists of labels $y$ flipped with a certain probability $\eta$ without depending on $\tilde{x}$.
In the proof of Proposition 28, we use the following lemmas.
Lemma 29 (Corollary 4.2.13 in [32]).
The covering numbers of the Euclidean ball $B_{n}^{2}:=\{x\in\mathbb{R}^{n}\mid\|x\|\leq 1\}$ satisfy the following for any $\epsilon>0$:
$\left(\frac{1}{\epsilon}\right)^{n}\leq\mathcal{N}(B_{n}^{2},\epsilon)\leq\left(\frac{2}{\epsilon}+1\right)^{n}.$
The same upper bound holds for the unit Euclidean sphere $S^{n-1}$.
Lemma 30 (Exercise 4.4.3 in [32]).
Let $A$ be an $n\times n$ real symmetric matrix and $\epsilon\in[0,1/2)$. For any $\epsilon$-net $\mathcal{N}_{\epsilon}$ of the sphere $S^{n-1}$,
$\underset{x\in\mathcal{N}_{\epsilon}}{\mathrm{sup}}|(Ax)\cdot x|\leq\|A\|_{\mathrm{op}}\leq\frac{1}{1-2\epsilon}\underset{x\in\mathcal{N}_{\epsilon}}{\mathrm{sup}}|(Ax)\cdot x|.$
Lemma 31 (Lemma A.3 in [13]).
For any $\alpha\in(0,1)$ and any random variables $X,Y$ we have
$\|X+Y\|_{\psi_{\alpha}}\leq 2^{1/\alpha}(\|X\|_{\psi_{\alpha}}+\|Y\|_{\psi_{\alpha}}).$
Lemma 32 (Lemma 4.1.5 in [32]).
Let $A$ be an $m\times n$ real matrix and $\delta>0$. Suppose that
$\|A^{T}A-I_{n}\|_{\mathrm{op}}\leq\max(\epsilon,\epsilon^{2}).$
Then
$(1-\epsilon)\|x\|\leq\|Ax\|\leq(1+\epsilon)\|x\|\quad\text{for all }x\in\mathbb{R}^{n}.$
Consequently,
$1-\epsilon\leq s_{k}(A)\leq 1+\epsilon\quad\text{for all }k\in[n].$
Proof of Proposition 28.
By Lemma 29, there exists a $\frac{1}{4}$-net $\mathcal{N}_{1/4}$ of the unit sphere $S^{n-1}$ with cardinality $|\mathcal{N}_{1/4}|\leq 9^{n}$. By Lemma 30, we have
$\left\|\frac{1}{\sigma^{2}p}\tilde{X}^{T}\tilde{X}-I_{n}\right\|_{\mathrm{op}}$ $\leq 2\underset{u\in\mathcal{N}_{1/4}}{\max}\left|\left(\left(\frac{1}{\sigma^{2}p}\tilde{X}^{T}\tilde{X}-I_{n}\right)u\right)\cdot u\right|$
$=2\underset{u\in\mathcal{N}_{1/4}}{\max}\left|\frac{1}{\sigma^{2}p}\|\tilde{X}u\|^{2}-1\right|.$ (7)
Fix $u\in S^{n-1}$. Let $r_{i}\in\mathbb{R}^{n}$ denote the $i$-th row of $Q=[q_{1},\cdots,q_{n}]\in\mathbb{R}^{p\times n}$. We have
$\frac{1}{\sigma^{2}p}\|\tilde{X}u\|^{2}$
$=\frac{1}{p}\sum_{i=1}^{p}\frac{1}{\sigma^{2}}((r_{i}\odot\tilde{y}+\mu_{i}\mathbf{1})\cdot u)^{2}$
$=\frac{1}{p}\sum_{i=1}^{p}\frac{1}{\sigma^{2}}\left(((r_{i}\odot\tilde{y})\cdot u)^{2}+2((r_{i}\odot\tilde{y})\cdot u)\sum_{j=1}^{n}\mu_{i}u_{j}+\mu_{i}^{2}\left(\sum_{j=1}^{n}u_{j}\right)^{2}\right)$
$=\frac{1}{p}\sum_{i=1}^{p}\left(f_{i}+\frac{\mu_{i}^{2}}{\sigma^{2}}\left(\sum_{j=1}^{n}u_{j}\right)^{2}\right),$
where $f_{i}=\frac{1}{\sigma^{2}}\left((r_{i}\cdot(\tilde{y}\odot u))^{2}+2(r_{i}\cdot(\tilde{y}\odot u))\mu_{i}\sum_{j=1}^{n}u_{j}\right)$. Thus,
$\mathbb{P}\left[\left|\frac{1}{\sigma^{2}p}\|Xu\|^{2}-1\right|>\frac{\epsilon}{2}\right]$
$=\mathbb{P}\left[\left|\frac{1}{p}\sum_{i=1}^{p}\left(f_{i}+\frac{\mu_{i}^{2}}{\sigma^{2}}\left(\sum_{j=1}^{n}u_{j}\right)^{2}\right)-1\right|>\frac{\epsilon}{2}\right]$
$=\sum_{\tilde{y}\in\{-1,1\}^{n}}\underset{Q\sim P_{\mathrm{clust}}^{n}}{\mathbb{P}}\left[\left|\frac{1}{p}\sum_{i=1}^{p}\left(f_{i}+\frac{\mu_{i}^{2}}{\sigma^{2}}\left(\sum_{j=1}^{n}u_{j}\right)^{2}\right)-1\right|>\frac{\epsilon}{2}\right]2^{-n}$
$=\sum_{\tilde{y}\in\{-1,1\}^{n}}\underset{Q\sim P_{\mathrm{clust}}^{n}}{\mathbb{P}}\left[\frac{1}{p}\sum_{i=1}^{p}f_{i}-1>\frac{\epsilon}{2}-\frac{\|\mu\|^{2}}{\sigma^{2}p}\left(\sum_{j=1}^{n}u_{j}\right)^{2}\right]2^{-n}$
$\quad+\sum_{\tilde{y}\in\{-1,1\}^{n}}\underset{Q\sim P_{\mathrm{clust}}^{n}}{\mathbb{P}}\left[\frac{1}{p}\sum_{i=1}^{p}f_{i}-1<-\frac{\epsilon}{2}-\frac{\|\mu\|^{2}}{\sigma^{2}p}\left(\sum_{j=1}^{n}u_{j}\right)^{2}\right]2^{-n}.$
$\leq\sum_{\tilde{y}\in\{-1,1\}^{n}}\underset{Q\sim P_{\mathrm{clust}}^{n}}{\mathbb{P}}\left[\frac{1}{p}\sum_{i=1}^{p}f_{i}-1>\frac{\epsilon}{2}-\frac{n\|\mu\|^{2}}{\sigma^{2}p}\right]2^{-n}$
$\quad+\sum_{\tilde{y}\in\{-1,1\}^{n}}\underset{Q\sim P_{\mathrm{clust}}^{n}}{\mathbb{P}}\left[\frac{1}{p}\sum_{i=1}^{p}f_{i}-1<-\frac{\epsilon}{2}\right]2^{-n}.$
The final inequality was derived using $\sum_{j=1}^{n}u_{j}\leq\sqrt{n}$. $\underset{Q\sim P_{\mathrm{clust}}^{n}}{\mathbb{E}}[f_{i}]=1$ and $\{f_{i}\}_{i=1}^{p}$ are independent random variables when conditioned on $\tilde{y}$. All elements of $Q$ are $\alpha$ sub-exponential with their exponential Orlicz norm at most $1$. By Proposition 18, there is a constant $c$ depending on $\alpha$ such that for any $t\geq 0$,
$\mathbb{P}[|r_{i}\cdot(\tilde{y}\odot u)|\geq t]$ $\leq 2\exp\left(-c\frac{t^{\alpha}}{\|\tilde{y}\odot u\|^{\alpha}}\right)=2\exp\left(-ct^{\alpha}\right).$
Thus, there is a constant $K_{1}$ such that
$\|r_{i}\cdot(\tilde{y}\odot u)\|_{\psi_{\alpha}}\leq K_{1}.$
Then, we have
$\|(r_{i}\cdot(\tilde{y}\odot u))^{2}\|_{\psi_{\alpha/2}}$ $=\inf\left\{t>0:\mathbb{E}\left[\exp\left(\frac{|r_{i}\cdot(\tilde{y}\odot u)|^{\alpha}}{\sqrt{t}^{\alpha}}\right)\right]\leq 2\right\}\leq\sqrt{K_{1}},$
and there is a constant $K_{2}$ such that
$\|r_{i}\cdot(\tilde{y}\odot u)\|_{\psi_{\alpha/2}}\leq K_{2}.$
By Lemma 31 and the fact that $\|\cdot\|_{\psi_{1}}$ preserves the triangle inequality, we obtain
$\|f_{i}\|_{\psi_{\alpha/2}}$ $=\frac{1}{\sigma^{2}}\left\|(r_{i}\cdot(\tilde{y}\odot u))^{2}+2(r_{i}\cdot(\tilde{y}\odot u))\mu_{i}\sum_{j=1}^{n}u_{j}\right\|_{\psi_{\alpha/2}}$
$\leq\frac{2^{2/\alpha}}{\sigma^{2}}\left(\left\|(r_{i}\cdot(\tilde{y}\odot u))^{2}\right\|_{\psi_{\alpha/2}}+\left\|2(r_{i}\cdot(\tilde{y}\odot u))\mu_{i}\sum_{j=1}^{n}u_{j}\right\|_{\psi_{\alpha/2}}\right)$
$=\frac{2^{2/\alpha}}{\sigma^{2}}\left(\sqrt{K_{1}}+\left|2\mu_{i}\ \sum_{j=1}^{n}u_{j}\right|K_{2}\right)$
$\leq\frac{2^{2/\alpha}}{\sigma^{2}}\left(\sqrt{K_{1}}+2\max_{i}|\mu_{i}|\sqrt{n}K_{2}\right).$
Setting $\frac{\epsilon}{2}\geq\frac{n\|\mu\|^{2}}{\sigma^{2}p}$, by Proposition 18, there exists a constant $c$ such that
$\mathbb{P}\left[\left|\frac{1}{\sigma^{2}p}\|\tilde{X}u\|^{2}-1\right|>\frac{\epsilon}{2}\right]$ $\leq\sum_{\tilde{y}\in\{-1,1\}^{n}}2\exp\left(-\frac{\sigma^{\alpha}}{2c}\frac{\left(\frac{\epsilon}{2}-\frac{n\|\mu\|^{2}}{\sigma^{2}p}\right)^{\alpha/2}p^{\alpha/2}}{\left(\sqrt{K_{1}}+2\max_{i}|\mu_{i}|\sqrt{n}K_{2}\right)^{\alpha/2}}\right)2^{-n}$
$\quad+\sum_{\tilde{y}\in\{-1,1\}^{n}}2\exp\left(-\frac{\sigma^{\alpha}}{2c}\frac{\left(\frac{\epsilon}{2}\right)^{\alpha/2}p^{\alpha/2}}{\left(\sqrt{K_{1}}+2\max_{i}|\mu_{i}|\sqrt{n}K_{2}\right)^{\alpha/2}}\right)2^{-n}.$
By setting
$\frac{\epsilon}{2}=\frac{n\|\mu\|^{2}}{\sigma^{2}p}+\frac{(2c)^{2/\alpha}\left(\sqrt{K_{1}}+2\max_{i}|\mu_{i}|\sqrt{n}K_{2}\right)}{\sigma^{2}p}\left(n\log 9+\log\frac{4}{\delta}\right)^{\frac{2}{\alpha}},$
we have
$\mathbb{P}\left[\left|\frac{1}{\sigma^{2}p}\|\tilde{X}u\|^{2}-1\right|>\frac{\epsilon}{2}\right]$ $\leq\sum_{\tilde{y}\in\{-1,1\}^{n}}4\exp\left(-n\log 9-\log\frac{4}{\delta}\right)2^{-n}$
$=4\exp\left(-n\log 9-\log\frac{4}{\delta}\right).$
By inequality (7) and the union bound method, we have
$\left\|\frac{1}{\sigma^{2}p}\tilde{X}^{T}\tilde{X}-I_{n}\right\|_{\mathrm{op}}$ $\leq 2\underset{u\in\mathcal{N}_{1/4}}{\max}\left|\left(\left(\frac{1}{\sigma^{2}p}\tilde{X}^{T}\tilde{X}-I_{n}\right)u\right)\cdot u\right|$
$\leq\mathbb{P}\left[\underset{u\in\mathcal{N}_{1/4}}{\max}\left|\frac{1}{\sigma^{2}p}\|\tilde{X}u\|^{2}-1\right|>\frac{\epsilon}{2}\right]$
$\leq 9^{n}\cdot 4\exp\left(-n\log 9-\log\frac{4}{\delta}\right)$
$=\delta.$
By Lemma 32, we conclude that
$s_{1}(\tilde{X})$
$\leq\sigma\sqrt{p}\left(1+\frac{2n\|\mu\|^{2}}{\sigma^{2}p}+\frac{2(2c)^{2/\alpha}\left(\sqrt{K_{1}}+2\max_{i}|\mu_{i}|\sqrt{n}K_{2}\right)}{\sigma^{2}p}\left(n\log 9+\log\frac{4}{\delta}\right)^{2/\alpha}\right),$
which completes our proof. A similar argument holds for $X=[y_{1}\tilde{x_{1}},\cdots,y_{n}\tilde{x_{n}}]$, which consists of labels $y$ flipped with a certain probability $\eta$ without depending on $\tilde{x}$. ∎
We will prove Proposition 6, which provides an upper bound for the singular values of $X$, by making a slight modification to the proof of Proposition 28. In the proof of Proposition 6 , we use the following lemma.
Lemma 33 (Lemma A.2 in [13] and Proposition 8.1 in [4]).
Let $d_{\alpha}:=\frac{(\alpha e)^{1/\alpha}}{2}$ and $D_{\alpha}:=(2e)^{1/\alpha}$ for $\alpha\in(0,1)$, and let $d_{\alpha}:=\frac{1}{2}$ and $D_{\alpha}:=2e$ for $\alpha\geq 1$. For any $\alpha>0$ and any random variable $X$, we have
$d_{\alpha}\sup_{p\geq 1}\|X\|_{L_{p}}\leq\|X\|_{\psi_{\alpha}}\leq D_{\alpha}\sup_{p\geq 1}\|X\|_{L_{p}}.$
Proof of Proposition 6.
Fix $u\in S^{n-1}$. Let $r_{i}\in\mathbb{R}^{n}$ denote the $i$-th row of $Q=[q_{1},\cdots,q_{n}]\in\mathbb{R}^{p\times n}$. We have
$\frac{1}{p}\|Xu\|^{2}$
$=\frac{1}{p}\sum_{i=1}^{p}((r_{i}\odot y+\mu_{i}y\odot\tilde{y})\cdot u)^{2}$
$=\frac{1}{p}\sum_{i=1}^{p}\left(((r_{i}\odot y)\cdot u)^{2}+2\mu_{i}((r_{i}\odot y)\cdot u)((y\odot\tilde{y})\cdot u)+\mu_{i}^{2}\left(\sum_{j=1}^{n}y_{j}\tilde{y}_{j}u_{j}\right)^{2}\right)$
$=\frac{1}{p}\sum_{i=1}^{p}\left(f_{i}+\mu_{i}^{2}\left(\sum_{j=1}^{n}y_{j}\tilde{y}_{j}u_{j}\right)^{2}\right),$
where $f_{i}=((r_{i}\odot y)\cdot u)^{2}+2\mu_{i}((r_{i}\odot y)\cdot u)((y\odot\tilde{y})\cdot u)$. Since $y\in\{-1,1\}$, each component of $r_{i}\odot y$ is $\alpha$ sub-exponential with their exponential Orlicz norm at most $1$. By Proposition 18, there is a constant $c$ depending on $\alpha$ such that for any $t\geq 0$,
$\mathbb{P}[|(r_{i}\odot y)\cdot u|\geq t]$ $\leq 2\exp\left(-c\frac{t^{\alpha}}{\|u\|^{\alpha}}\right)=2\exp\left(-ct^{\alpha}\right).$
Thus, there is a constant $K_{1}$ such that
$\|(r_{i}\odot y)\cdot u\|_{\psi_{\alpha}}\leq K_{1},$
and
$\|2\mu_{i}((r_{i}\odot y)\cdot u)((y\odot\tilde{y})\cdot u)\|_{\psi_{\alpha}}$ $=2|\mu_{i}|\left\|((r_{i}\odot y)\cdot u)((y\odot\tilde{y})\cdot u)\right\|_{\psi_{\alpha}}$
$\leq 2|\mu_{i}|\left\||(r_{i}\odot y)\cdot u||(y\odot\tilde{y})\cdot u|\right\|_{\psi_{\alpha}}$
$\leq 2|\mu_{i}|\sum_{j=1}^{n}|u_{j}|\left\|(r_{i}\odot y)\cdot u\right\|_{\psi_{\alpha}}$
$\leq 2K_{1}|\mu_{i}|\sum_{j=1}^{n}|u_{j}|.$
By Lemma 33, there is a constant $K_{2}$ and $K_{3}$ such that for any $p\geq 1$,
$\|(r_{i}\odot y)\cdot u\|_{L_{p}}\leq K_{3}p^{1/\alpha},$
$\|2\mu_{i}((r_{i}\odot y)\cdot u)((y\odot\tilde{y})\cdot u)\|_{L^{p}}\leq K_{2}|\mu_{i}|\sum_{j=1}^{n}|u_{j}|p^{1/\alpha},$
Therefore,
$\mathbb{E}[((r_{i}\odot y)\cdot u)^{2}]\in[0,2^{1/\alpha}K_{3}],$
$\mathbb{E}[2\mu_{i}((r_{i}\odot y)\cdot u)((y\odot\tilde{y})\cdot u)]\in\left[-K_{2}|\mu_{i}|\sum_{j=1}^{n}|u_{j}|,K_{2}|\mu_{i}|\sum_{j=1}^{n}|u_{j}|\right].$
By combining these two, we have
$\mathbb{E}[f_{i}]\in\left[-K_{2}|\mu_{i}|\sum_{j=1}^{n}|u_{j}|,2^{1/\alpha}K_{3}+K_{2}|\mu_{i}|\sum_{j=1}^{n}|u_{j}|\right].$
Thus,
$\mathbb{P}\left[\left|\frac{1}{p}\|Xu\|^{2}-1\right|>\frac{\epsilon}{2}\right]$
$=\mathbb{P}\left[\left|\frac{1}{p}\sum_{i=1}^{p}\left(f_{i}+\mu_{i}^{2}\left(\sum_{j=1}^{n}y_{j}\tilde{y}_{j}u_{j}\right)^{2}\right)-1\right|>\frac{\epsilon}{2}\right]$
$=\mathbb{P}\left[\frac{1}{p}\sum_{i=1}^{p}(f_{i}-\mathbb{E}[f_{i}])>\frac{\epsilon}{2}+1-\frac{1}{p}\sum_{i=1}^{p}\mathbb{E}[f_{i}]-\frac{\|\mu\|^{2}}{p}\left(\sum_{j=1}^{n}y_{j}\tilde{y}_{j}u_{j}\right)^{2}\right]$
$\quad+\mathbb{P}\left[\frac{1}{p}\sum_{i=1}^{p}(f_{i}-\mathbb{E}[f_{i}])<-\frac{\epsilon}{2}+1-\frac{1}{p}\sum_{i=1}^{p}\mathbb{E}[f_{i}]-\frac{\|\mu\|^{2}}{p}\left(\sum_{j=1}^{n}y_{j}\tilde{y}_{j}u_{j}\right)^{2}\right]$
$\leq\mathbb{P}\left[\frac{1}{p}\sum_{i=1}^{p}(f_{i}-\mathbb{E}[f_{i}])>\frac{\epsilon}{2}-\left(2^{1/\alpha}K_{3}+\frac{K_{2}}{p}\sum_{i=1}^{p}|\mu_{i}|\sum_{j=1}^{n}|u_{j}|\right)-\frac{n\|\mu\|^{2}}{p}\right]$
$\quad+\mathbb{P}\left[\frac{1}{p}\sum_{i=1}^{p}(f_{i}-\mathbb{E}[f_{i}])<-\frac{\epsilon}{2}+1-\left(-\frac{K_{2}}{p}\sum_{i=1}^{p}|\mu_{i}|\sum_{j=1}^{n}|u_{j}|\right)\right]$
$\leq\mathbb{P}\left[\frac{1}{p}\sum_{i=1}^{p}(f_{i}-\mathbb{E}[f_{i}])>\frac{\epsilon}{2}-\left(2^{1/\alpha}K_{3}+\frac{K_{2}}{p}\sqrt{n}\sum_{i=1}^{p}|\mu_{i}|\right)-\frac{n\|\mu\|^{2}}{p}\right]$
$\quad+\mathbb{P}\left[\frac{1}{p}\sum_{i=1}^{p}(f_{i}-\mathbb{E}[f_{i}])<-\frac{\epsilon}{2}+1+\frac{K_{2}}{p}\sqrt{n}\sum_{i=1}^{p}|\mu_{i}|\right].$
By the same method as the proof of Proposition 28, we have
$\|f_{i}\|_{\psi_{\alpha/2}}\leq 2^{2/\alpha}\left(\sqrt{K_{1}}+2\max_{i}|\mu_{i}|\sqrt{n}K_{2}\right).$
Setting $\frac{\epsilon}{2}\geq 2^{1/\alpha}K_{3}+\frac{K_{2}}{p}\sqrt{n}\sum_{i=1}^{p}|\mu_{i}|+\frac{n\|\mu\|^{2}}{p}+1$, by Proposition 18, there exists a constant $c$ such that
$\mathbb{P}\left[\left|\frac{1}{p}\|\tilde{X}u\|^{2}-1\right|>\frac{\epsilon}{2}\right]$
$\leq\sum_{\tilde{y}\in\{-1,1\}^{n}}2\exp\left(-\frac{1}{2c}\frac{\left(\frac{\epsilon}{2}-\left(2^{1/\alpha}K_{3}+\frac{K_{2}}{p}\sqrt{n}\sum_{i=1}^{p}|\mu_{i}|\right)-\frac{n\|\mu\|^{2}}{p}\right)^{\alpha/2}p^{\alpha/2}}{\left(\sqrt{K_{1}}+2\max_{i}|\mu_{i}|\sqrt{n}K_{2}\right)^{\alpha/2}}\right)2^{-n}$
$\quad+\sum_{\tilde{y}\in\{-1,1\}^{n}}2\exp\left(-\frac{1}{2c}\frac{\left(\frac{\epsilon}{2}-1-\frac{K_{2}}{p}\sqrt{n}\sum_{i=1}^{p}|\mu_{i}|\right)^{\alpha/2}p^{\alpha/2}}{\left(\sqrt{K_{1}}+2\max_{i}|\mu_{i}|\sqrt{n}K_{2}\right)^{\alpha/2}}\right)2^{-n}.$
By setting
$\frac{\epsilon}{2}$ $=2^{1/\alpha}K_{3}+\frac{K_{2}}{p}\sqrt{n}\sum_{i=1}^{p}|\mu_{i}|+\frac{n\|\mu\|^{2}}{p}+1$
$\quad+\frac{(2c)^{2/\alpha}\left(\sqrt{K_{1}}+2\max_{i}|\mu_{i}|\sqrt{n}K_{2}\right)}{p}\left(n\log 9+\log\frac{4}{\delta}\right)^{\frac{2}{\alpha}},$
we have
$\mathbb{P}\left[\left|\frac{1}{p}\|Xu\|^{2}-1\right|>\frac{\epsilon}{2}\right]$ $\leq 4\exp\left(-n\log 9-\log\frac{4}{\delta}\right).$
The remainder of the proof is the same as the proof of Proposition 28. ∎
B.3 Proofs of Corollary 8 and 9
Proof of Corollary 8 and 9.
We have
$\sum_{i=1}^{p}|\mu_{i}|$ $\leq\sqrt{p}\|\mu\|,$
and
$\max_{i}|\mu_{i}|\leq\|\mu\|.$
Under assumptions (A3) and (A4), we have the following inequalities:
  • $p\geq C\|\mu\|^{2}n$.
  • $p\geq C\|\mu\|n^{3/2}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}$.
  • $p\geq Cn^{2}\left(\log\frac{n}{\delta}\right)^{\frac{2}{\alpha}}$.
Therefore, we have
$c_{5}+\frac{c_{6}\sqrt{n}}{p}\sum_{i=1}^{p}|\mu_{i}|+\frac{2n\|\mu\|^{2}}{p}$
$\quad+\frac{c_{7}+c_{8}\max_{i}|\mu_{i}|\sqrt{n}}{p}\left(n\log 9+\log\frac{4}{\delta}\right)^{\frac{2}{\alpha}}$
$\leq c_{5}+\frac{c_{6}\sqrt{n}\|\mu\|}{\sqrt{p}}+\frac{2n\|\mu\|^{2}}{p}$
$\quad+\frac{c_{7}+c_{8}\|\mu\|\sqrt{n}}{p}\left(n\log 9+\log\frac{4}{\delta}\right)^{\frac{2}{\alpha}}$
$\leq c_{5}+\frac{c_{6}}{C}n^{-1}\left(\log\frac{n}{\delta}\right)^{-\frac{1}{\alpha}}+\frac{2}{C}$
$\quad+\frac{c_{7}}{C}n^{-2}\left(\log\frac{n}{\delta}\right)^{-\frac{2}{\alpha}}+\frac{c_{8}}{C}n^{-1}\left(\log\frac{n}{\delta}\right)^{-\frac{1}{\alpha}}\left(n\log 9+\log\frac{4}{\delta}\right)^{\frac{2}{\alpha}}$
$=O\left(1+n^{\frac{2}{\alpha}-1}(\log n)^{-\frac{1}{\alpha}}\right).$
By performing a similar calculation, we obtain
$1+\frac{2n}{p}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)=O(1).$
Therefore, if we regard $n$ as fixed, we have
$\min\Bigg{(}\frac{8}{p}$ $\Bigg{(}c_{5}+\frac{c_{6}\sqrt{n}}{p}\sum_{i=1}^{p}|\mu_{i}|+\frac{2n\|\mu\|^{2}}{p}$
$\quad+\frac{c_{7}+c_{8}\max_{i}|\mu_{i}|\sqrt{n}}{p}\left(n\log 9+\log\frac{4}{\delta}\right)^{\frac{2}{\alpha}}\Bigg{)}^{-2},$
$\frac{1}{c_{2}p}\left(1+\frac{2n}{p}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right)^{-1}\Bigg{)}$
$=O(p).$
On the other hand, if we regard $n$ as not fixed, we have
$\min\Bigg{(}\frac{8}{p}$ $\Bigg{(}c_{5}+\frac{c_{6}\sqrt{n}}{p}\sum_{i=1}^{p}|\mu_{i}|+\frac{2n\|\mu\|^{2}}{p}$
$\quad+\frac{c_{7}+c_{8}\max_{i}|\mu_{i}|\sqrt{n}}{p}\left(n\log 9+\log\frac{4}{\delta}\right)^{\frac{2}{\alpha}}\Bigg{)}^{-2},$
$\frac{1}{c_{2}p}\left(1+\frac{2n}{p}\left(\|\mu\|^{2}+\sqrt{p}\left(\log\frac{n}{\delta}\right)^{\frac{1}{\alpha}}\right)\right)^{-1}\Bigg{)}$
$=O\left(p^{-1}\left(1+n^{\frac{2}{\alpha}-1}(\log n)^{-\frac{1}{\alpha}}\right)^{-2}\right).$
Appendix C Details of the Figures in the introduction
C.1 Figure 1 : An example of input in image analysis exhibiting heavier tails than sub-gaussian
In this section, we demonstrate that some of the feature representations derived from real-world image datasets exhibit distributions heavier than sub-gaussian distributions.
C.1.1 Methodology
To estimate the tail-heaviness of feature distributions, we followed the steps below:
  1. 1.
    A total of $n$ samples were collected from intermediate layers of CNN models trained on several image datasets.
  2. 2.
    Each sample of feature value was centered by subtracting its mean, and the absolute value of the result was taken.
  3. 3.
    The upper 5% of the sorted absolute values was selected, yielding order statistics $(x^{(1)},x^{(2)},\dots,x^{(\lfloor 0.05n\rfloor)})$.
  4. 4.
    For each $x^{(i)}$, we computed the corresponding $z_{i}=-\log(i/n)$, which approximates $-\log\mathbb{P}[|X|\geq x^{(i)}]$.
  5. 5.
    We then performed a regression of $(x^{(i)},z_{i})$ against the form $f(t)=at^{\xi}+b$, enabling us to estimate the tail parameter $\xi$. The regression was performed using the non-linear least squares method implemented via the scipy.optimize.curve_fit function.
The tail parameter $\xi$ is crucial as it characterizes the heaviness of the distribution’s tail, where $P(|X|\geq t)=\exp(-(at^{\xi}+b))$.
Table 1: Summary of datasets used in the simulation.
Dataset Split Number of Images Image Size Number of Labels
CIFAR-10 [17] Training 50,000 32x32 10
Test 10,000 32x32 10
CIFAR-100 [17] Training 50,000 32x32 100
Test 10,000 32x32 100
Fashion-MNIST [36] Training 60,000 28x28 10
Test 10,000 28x28 10
SVHN [28] Training 73,257 32x32 10
Test 26,032 32x32 10
Table 2: Mean and Variance of Estimated Tail Index ($\xi$)
Dataset Mean ($\hat{\xi}$) Variance ($\hat{\xi}$)
CIFAR-10 0.9771 0.1111
CIFAR-100 1.0423 0.1281
Fashion-MNIST 1.4748 0.7295
SVHN 0.9272 0.0514
Gaussian 1.6172 0.2388
Exponential 0.8996 0.0535
Figure 6: Histograms of estimated tail index ($\xi$)
C.1.2 Datasets and Models
For the simulations, we used intermediate layer outputs from CNN models trained on the datasets listed in Table 1.
The CNN architecture used for these datasets consisted of three convolutional layers (with 32, 64, and 64 filters, respectively) followed by max-pooling layers. The final fully connected layers had 64 neurons, with ReLU activations throughout the network. The total number of parameters in the network is typical for small-scale models. The output layer size was adjusted according to the number of classes in each dataset (e.g., 100 for CIFAR-100). This CNN model was trained using the Adam optimizer with a learning rate of 0.001 and cross-entropy loss as the loss function. Training was performed over 100 epochs, with a batch size of 100. All images were normalized and converted to PyTorch tensors prior to training.
The CNN models were trained using the training split of the datasets listed in Table 1. After training, the test data from each dataset was used to generate intermediate layer outputs, which were then used in our analysis to evaluate the feature distributions.
To compare these real-world results, we also generated samples from Gaussian and exponential distributions as baseline comparisons, aligning the future vector size with the smallest intermediate output size in this simulation, which is $3136$, and matching the sample size to the smaller value of $10000$.
C.1.3 Results
Table 2 and Figures 1 and 6 revealed that some intermediate layer outputs from CNN models exhibit distributions with heavier tails than Gaussian distributions. This finding reveals the necessity of developing a theory that addresses heavy-tailed distributions.
It should be noted that, in the case of the Gaussian distribution, the reason why the value of $\xi$ is distributed below $2$ is that, in the samples used for the calculation, the approximation $P(|X|>t)\simeq 2\exp(-t^{2}/2)$ is not sufficiently accurate. When limited to the samples further in the tail of the distribution, the values of $\xi$ approach $2$.
C.2 Figure 2 : Benign overfitting can occur even for heavy inputs
In this section, we conduct simulations to demonstrate that benign overfitting can occur even in settings with input distributions heavier-tailed than sub-gaussian. Specifically, we analyze the performance of a linear classifier trained using gradient descent on data drawn from generalized normal distributions, investigating the relationship between dimensionality $p$, tail heaviness (controlled by the shape parameter $\gamma$), and the classifier’s error rates.
C.2.1 Data Generation
Table 3: Test Error Data with Mean and Standard Error of the Mean for Different $\gamma$ Values
Dimension (p) $\gamma$ Mean Test Error Test Error SEM
100 0.25 0.1288 0.0047
0.5 0.1047 0.0049
2 0.0823 0.0040
200 0.25 0.0808 0.0035
0.5 0.0660 0.0020
2 0.0627 0.0021
300 0.25 0.0652 0.0018
0.5 0.0562 0.0014
2 0.0531 0.0012
400 0.25 0.0578 0.0015
0.5 0.0519 0.0012
2 0.0531 0.0009
500 0.25 0.0556 0.0013
0.5 0.0510 0.0009
2 0.0514 0.0008
600 0.25 0.0525 0.0010
0.5 0.0493 0.0009
2 0.0504 0.0011
700 0.25 0.0513 0.0007
0.5 0.0483 0.0009
2 0.0515 0.0012
800 0.25 0.0518 0.0011
0.5 0.0495 0.0011
2 0.0512 0.0012
900 0.25 0.0502 0.0009
0.5 0.0503 0.0010
2 0.0500 0.0009
1000 0.25 0.0501 0.0010
0.5 0.0508 0.0009
2 0.0508 0.0009
1100 0.25 0.0526 0.0011
0.5 0.0511 0.0009
2 0.0512 0.0011
1200 0.25 0.0518 0.0009
0.5 0.0499 0.0007
2 0.0507 0.0010
1300 0.25 0.0499 0.0010
0.5 0.0498 0.0010
2 0.0498 0.0008
1400 0.25 0.0492 0.0011
0.5 0.0499 0.0008
2 0.0492 0.0009
1500 0.25 0.0502 0.0010
0.5 0.0500 0.0010
2 0.0497 0.0010
The data was generated under the heavy-tailed setting, as described in Section 2.3. The specific configuration is as follows:
  • $p$ : We varied the number of features $p$ from 100 to 1500 in increments of 100, to study the effect of increasing dimensionality on the model’s performance.
  • $n_{\text{train}}$ and $n_{\text{test}}$ : For the training data, we used $n=200$ samples, while the test data consisted of $n_{\text{test}}=1000$ samples.
  • $P_{\mathrm{clust}}$: Each component of $P_{\mathrm{clust}}$ is independently and identically distributed according to a generalized normal distribution, with specified location, scale, and shape parameters.
    • The location parameter is $\mathbf{0}$.
    • The shape parameters $\gamma$ are $0.25$, $0.5$, and $2$
    • The scale parameter $\sigma$ is adjusted for each $\gamma$ such that the variance is fixed at $1$.
  • $\mu$ : The mean vector $\mu$ was set as $\mu=\mathbf{1}$, meaning that all features had a common shift.
  • $U$ : We applied an orthogonal transformation to the samples using a matrix $U$, which was obtained from the QR decomposition of a randomly generated matrix $A$. Each element of $A$ was drawn from a standard normal distribution. The orthogonal matrix $U$ is the result of the decomposition:
    $A=UR$
    where $R$ is an upper triangular matrix.
  • $\eta$ : For each sample, we generated a label $y\in\{-1,1\}$ by multiplying a random scalar by a noise factor $\eta$, where $\eta=0.05$ in all experiments.
C.2.2 Model training
We used the maximum margin classifier, as described in Section 2.4. The model was trained for $100000$ epochs to ensure convergence. The learning rate was set to $\beta=0.001$. Each experiment was repeated $50$ times, and the results were averaged. To ensure robustness, $95$% confidence intervals were calculated based on the standard error of the mean.
C.3 Results
From Figure 2 and 3, we can observe that the training error remains near zero across all dimensions, while the test error initially decreases and then stabilizes around the noise level as the dimension increases. This indicates that benign overfitting can occur even with distributions that have heavier tails than sub-gaussian distributions.
Appendix D Experimental code and computing infrastructure
The experimental code can be obtained from the anonymous URL on OSF:https://osf.io/g6n9u/?view_only=f37a41efbee4421e8aa877f48c5879b4
The experiments were conducted in the following infrastructure:
  • GPU Type: NVIDIA GeForce RTX 4090
  • Number of GPUs: Single GPU
  • CPU Specifications: 13th Gen Intel(R) Core(TM) i9-13900KF 3.00 GHz
  • Memory: 32.0 GB
  • Operating System: Windows 10 Home 23H2
  • Frameworks and Libraries: The experiments for other figures were run using PyTorch, NumPy, SciPy, Matplotlib, Seaborn and Pandas on this infrastructure.